### v1

# Unified Modeling and Structural Optimization of Multi-magnet Embedded Soft Continuum Robots for Enhanced Kinematic Performances

Abstract
This paper presents a unified modeling and optimization framework to enhance the kinematic performance of multi-magnet embedded soft continuum robots (MeSCRs). To this end, we establish a differentiable system formulation based on an extended pseudo-rigid-body model. This formulation enables analysis of the equilibrium well-posedness and the geometry of the induced configuration under magnetic actuation. In particular, we show that the maximum controllable degrees of freedom of a MeSCR equal twice the number of embedded magnets. We subsequently develop a structural optimization framework based on differential geometry that links classical kinematic measures (e.g., manipulability and dexterity) to the configuration of embedded magnets. The resulting optimization condition reveals that improving local performance requires structurally modulating the spectrum of the configuration space metric to counteract its distortion. Closed-form solutions for optimal magnet configurations are derived under representative conditions, and a gradient-based numerical method is proposed for general design scenarios. Simulation studies validate the effectiveness of the proposed framework.
Index Terms:
Magnetic soft continuum robots, pseudo-rigid-body modeling, kinematic performances, structure optimization.
I Introduction
Magnetically actuated soft continuum robots have demonstrated significant potential in medical procedures such as vascular intervention and gastrointestinal examination[1]. Early designs typically embedded a single magnet at the distal end to enable basic magnetic manipulation[2, 3, 4]. However, the limited torque generated by such configurations restricts the controllable degrees of freedom (DoF) less than two[5], significantly constraining spatial maneuverability. Recent advances in magnetic materials and microfabrication have enabled the integration of multiple micro-magnets along the continuum body[6, 7], giving rise to multi-magnet embedded soft continuum robots (MeSCRs). These systems provide multiple distributed sources of actuation torque, substantially enhancing kinematic performance and expanding the design space. Various strategies have been proposed to exploit these capabilities, such as using movable magnets to enlarge reachable workspace[8], designing magnet layouts for multimodal deformation[9, 10], or inducing axial twisting via circumferential arrangements[11]. These efforts can be seen as structural design choices, analogous to tuning joint orientations or link lengths in rigid manipulators to improve kinematic capabilities.
Despite this progress, the absence of a unified modeling framework for MeSCRs remains a major barrier. Unlike rigid-link manipulators, which benefit from well-established analytic tools for performance analysis and optimization, soft continuum robots lack a theoretical foundation that systematically links structure to kinematic function[12]. As a result, many mature strategies developed in rigid manipulators cannot be directly applied to MeSCRs, motivating the need for a new modeling paradigm that bridges this gap.
Current MeSCR modeling approaches typically fall into two categories: Cosserat rod models and pseudo-rigid-body (PRB) models. The Cosserat formulation provides high-fidelity representations and has been integrated with magnetic dipole actuation models for motion control[13, 14]. However, it requires solving boundary value problems derived from variational principles, making it analytically intractable and computationally intensive for iterative design. By contrast, the PRB model discretizes flexible structures into rigid links with elastic joints[15, 16], yielding a Lagrangian formulation analogous to rigid manipulators. This enables unified expressions for forward kinematics and simplifies equilibrium computation via unconstrained finite-dimensional optimization[17]. Early applications in single-magnet MeSCRs have demonstrated effective Jacobian-based control strategies[18, 19], and rank-based analysis has shown that such systems can control at most two DoF[5]. Recent work has extended the PRB formulation to multi-magnet settings[20, 10], preliminarily validating its capacity to simulate complex deformations.
Nevertheless, the PRB model also suffers from several inherent limitations. The choice of joint discretization entails a trade-off between modeling fidelity and computational cost. In particular, due to the lack of theoretical analysis on well-posedness, existing methods typically use numerical methods to analyze the multiplicity and jump phenomena of solutions[21]. Moreover, to our best knowledge, complete closed-form expressions for Jacobian (simplified expression can be found in [20, 19]) and Hessian matrices under multi-magnet configurations have not been established.
To address the limitations outlined above, the first part of this paper investigates an extended-pseudo-rigid-body (EPRB) model[17, 19], which introduces serial spherical joints to enable unified modeling of spatial bending and torsional responses. The EPRB model possesses a generalized expressive capacity: when the elastic modulus approaches zero, it naturally degenerates into a classical rigid-link manipulator model. Conversely, as the number of embedded magnets tends to infinity, it approximates the behavior of a magnetic continuum[22]. Based on this formulation, we conduct a well-posedness analysis of the static equilibrium equations, encompassing uniqueness, boundedness, and convergence properties. Furthermore, we derive complete analytical expressions for the Jacobian and Hessian matrices under multi-magnet configurations, thereby providing theoretical support for both kinematic analysis and controller design.
Building upon the proposed EPRB model, the second part of this paper further demonstrates its ability to facilitate geometric analysis and structural optimization of soft continuum robots. We begin by rigorously proving a key structural result: the maximum controllable degrees of freedom in a magnetically actuated soft continuum robot is exactly twice the number of embedded magnets. This finding corroborates prior empirical observations reported in many existing literature[5, 20]. Additionally, we show that when magnets are arranged axially, the EPRB model naturally reduces to the classical PRB formulation, and a closed-form solution to the equilibrium configuration can be obtained under small-angle premises. As structural optimization often aims to enhance a robot’s kinematic performance, we draw inspiration from rigid manipulator design, where Jacobian-based kinematic measures (e.g., manipulability and dexterity) are widely employed to guide structural synthesis and task planning[23, 24]. We develop a differential geometry-based optimization framework around kinematic measures, which connects kinematic performances to the configuration of embedded magnets in MeSCR. Unlike rigid-link manipulators, whose structural and performance optimization can be directly formulated in joint space, soft continuum robots exhibit implicitly coupled configurations governed by elastomagnetic equilibrium. As such, optimization must be performed over the actuation or design space, with performance indices reparametrized through the configuration manifold. To resolve this, we reinterpret the equilibrium set as a smooth submanifold and apply tools from geometric measure theory to reparametrize global performance objectives over the actuation space. We derive closed-form optimal magnet configuration in representative cases and develop an efficient gradient-based numerical method for the general scenario.
Throughout the technical description, vectors and matrices are defined using lowercase and uppercase bold symbols, respectively, whereas scalars are not bolded.
II Modeling
This section introduces the extended pseudo-rigid-body model to describe the MeSCR. Complete closed-form expressions for the manipulator Jacobian and Hessian matrices are derived. We further investigate the well-posedness of the equilibrium solution. These results provide the necessary mathematical foundation for subsequent kinematic analysis and performance optimization.
II-A Forward Kinematics and Manipulator Jacobian
Figure 1: Illustration of the basic concept of the EPRB model on describing the MeSCR. The continuous body is discretized by a series of rotational rigid rods with joints and frames attached. The bending and twisting deformations of the MeSCR are described by three joint variables $\theta_{\mathbf{t}}^{i}$, $\theta_{\mathbf{u}}^{i}$, and $\theta_{\mathbf{v}}^{i}$.
To meet the demands of entering narrow spaces and achieving complex deformations, MeSCRs are typically designed with slender structures. Certain well-founded assumptions contribute to simplifying the model[25], such as the rod remaining free of shear deformation and its cross-section always being orthogonal to the centerline. Additionally, some classical PRB models[15] neglect both twisting and stretching of the continuum. In contrast, we assume only that the continuum is inextensible, with twisting characterized by extended variables [26]. Within the EPRB model framework, a continuum robot of length $L$ is described by a series of rigid rods. As shown in Fig. 1, $N$ spherical joints are distributed along the centerline of the MeSCR at points of interest, connected by $N$ rigid rods with lengths $l_{i},i=0,\dots,N-1$. We use the Voronoi region of the joint to scale the potential energies, which is defined as $\ell_{i}=(l_{i-1}+l_{i})/2,i=0,\dots,N-1$. An orthogonal local frame $\{\bar{\mathbf{t}}_{i},\bar{\mathbf{u}}_{i},\bar{\mathbf{v}}_{i}=\bar{\mathbf{t}}_{i}\times\bar{\mathbf{u}}_{i}\}$ is attached to each joint, with overbars indicating that the MeSCR is in a straightened reference state, and the tangent vector $\bar{\mathbf{t}}_{i}$ remains aligned with the direction of the successive rod. The local frame is used to determine the rotation axis for each joint, expressed as $\boldsymbol{\theta}_{i}=\theta_{\mathbf{t}}^{i}\bar{\mathbf{t}}+\theta_{\mathbf{u}}^{i}\bar{\mathbf{u}}+\theta_{\mathbf{v}}^{i}\bar{\mathbf{v}}\in\mathbb{R}^{3}$, where $\theta_{\mathbf{t}}^{i},\theta_{\mathbf{v}}^{i},\theta_{\mathbf{u}}^{i}\in\mathbb{R}$ are three independent joint variables. The associated rotation matrix is defined as $\mathbf{R}_{i}(\boldsymbol{\theta}_{i})=\mathbf{e}^{[\boldsymbol{\theta}_{i}]_{\times}}\in\mathrm{SO}(3)$, and we define a useful shorthand notation $(\mathbf{R}_{i}^{j})$ to represent the product of rotation matrices that $\mathbf{R}_{i}^{j}=\mathbf{R}_{i}\mathbf{R}_{i+1}\cdots\mathbf{R}_{j}$. With the notation, local frame axes in a deformed state (②), such as the tangent vector, can be easily computed by $\bar{\mathbf{t}}=\mathbf{R}_{0}^{i}\bar{\mathbf{t}}_{i}$ from the reference state (①). The resulting elastic strain on the $i$th joint will be used to compute the elastic potential energy later.
The configuration space $(\boldsymbol{\theta})$ of the MeSCR can be written as
$\boldsymbol{\theta}=\left[\boldsymbol{\theta}_{0},\dots,\boldsymbol{\theta}_{N-1}\right]^{\top}\in\mathbb{T}\subset\mathbb{R}^{3N_{m}}.$ (1)
Let a fixed base frame $\{s\}$ attach to joint 00, frame $\{b\}$ attach to the distal end, and ${}_{s}^{b}\bar{\mathbf{H}}\in\mathrm{SE}(3)$ be the associated distal pose when the MeSCR is in its zero position. The screw axes in the fixed base frame are of the form:
$\boldsymbol{\xi}_{i}=\begin{bmatrix}\boldsymbol{\theta}_{i}\\ -\boldsymbol{\theta}_{i}\times\bar{\mathbf{p}}_{i}\end{bmatrix}$
where $\bar{\mathbf{p}}_{i}$ is the reference position of the $i$th joint associated with $l_{i}$. Then, the forward kinematics of the MeSCR can be represented in the following product of exponentials:
${}_{s}^{b}\mathbf{H}(\boldsymbol{\theta})=\mathbf{e}^{\boldsymbol{\xi}_{0}^{\wedge}}\mathbf{e}^{\boldsymbol{\xi}_{1}^{\wedge}}\cdots\mathbf{e}^{\boldsymbol{\xi}_{N-1}^{\wedge}}{}_{s}^{b}\bar{\mathbf{H}}(\mathbf{0})\in\mathrm{SE}(3).$
We next present the analytical expression of the space Jacobian $\mathbf{J}_{\boldsymbol{\theta}}(\boldsymbol{\theta})\in\mathbb{R}^{6\times 3N}$, defined in fixed frame coordinates. The space Jacobian relates small changes in joint space to distal pose as $\delta\mathbf{h}=\mathbf{J}_{\boldsymbol{\theta}}\delta\boldsymbol{\theta}$, where $\mathbf{h}=\mathrm{Log}\left({}_{s}^{b}\mathbf{H}\right)\in\mathbb{R}^{6}$ represents the distal pose in Cartesian coordinates. To facilitate calculation, the Jacobian matrix is partitioned into blocks, with each block corresponding to a joint space group as
$\mathbf{J}_{\boldsymbol{\theta}}^{s}=\left[\frac{\partial\mathbf{h}}{\partial\boldsymbol{\theta}_{0}},\ \frac{\partial\mathbf{h}}{\partial\boldsymbol{\theta}_{1}},\ \cdots\ ,\frac{\partial\mathbf{h}}{\partial\boldsymbol{\theta}_{N-1}}\right]$ (2)
where the derivative with respect to $\boldsymbol{\theta}_{i}$ is given by
$\frac{\partial\mathbf{h}}{\partial\boldsymbol{\theta}_{i}}=\mathbf{Ad}_{\mathbf{e}^{\boldsymbol{\xi}_{0}^{\wedge}}\cdots\mathbf{e}^{\boldsymbol{\xi}_{i-1}^{\wedge}}}\mathbf{J}_{\mathbf{H}_{l}}^{\boldsymbol{\xi}_{i}}\frac{\partial\boldsymbol{\xi}_{i}}{\partial\boldsymbol{\theta}_{i}}\in\mathbb{R}^{6\times 3}$
with $\frac{\partial\boldsymbol{\xi}_{i}}{\partial\boldsymbol{\theta}_{i}}=\begin{bmatrix}\mathbf{I}_{3},\ -[\bar{\mathbf{p}}_{i}]_{\times}\end{bmatrix}^{\top}$, and $\mathbf{J}_{\mathbf{H}_{l}}^{\boldsymbol{\xi}_{i}}$ is the left Jacobian of the $\mathrm{SE}(3)$ manifold, as discussed in [27]. To conclude, we note the following relationship between the space Jacobian and the body Jacobian $(\mathbf{J}_{\boldsymbol{\theta}}^{b}(\boldsymbol{\theta}))$ in case of need:
$\mathbf{J}_{\boldsymbol{\theta}}^{b}=\mathbf{Ad}_{\mathbf{e}^{\boldsymbol{\xi}_{0}^{\wedge}}\cdots\mathbf{e}^{\boldsymbol{\xi}_{N-1}^{\wedge}}}^{-1}\mathbf{J}_{\boldsymbol{\theta}}^{s}.$ (3)
Also for brevity, $\mathbf{J}_{\boldsymbol{\theta}}$ will be referred to $\mathbf{J}_{\boldsymbol{\theta}}^{s}$ unless specified.
II-B Equations of Motion
Figure 2: Illustration of a quasi-static state of the MeSCR under an external magnetic field $\mathbf{b}(\mathbf{p}_{k})$. Multiple embedded micromagnets at corresponding joints generate magnetic torques $\mathbf{m}_{k}\times\mathbf{b}_{k}$ that induce deformation in the MeSCR. When the magnetic torque and elastic moments reach the equilibrium, the MeSCR attains the quasi-static state.
The MeSCR’s equations of motion are derived from the Lagrangian mechanics, which are formulated based on the MeSCR’s potential energy. As shown in Fig. 2, the MeSCR is allowed to bend, twist, and magnetic actuatable. Thus, the total potential energy features elastic strains and magnetic moments, which are decomposed into the elastic $(\mathcal{E}_{e})$ and magnetic $(\mathcal{E}_{m})$ potential energies as
$\mathcal{E}=\mathcal{E}_{e}+\mathcal{E}_{m}.$ (4)
The respective energies are formulated as follows.
II-B1 Elastic Potential Energy
In general deformation conditions, the elastic strain can be approximated as a linear function of the joint angle $\boldsymbol{\theta}_{i}$, [15, 28, 26]. The quadratic form of the elastic strain formulates the elastic potential energy as
$\mathcal{E}_{e}$ $=\frac{1}{2}\sum_{i=0}^{N-1}(\boldsymbol{\theta}_{i}-\bar{\boldsymbol{\theta}}_{i})^{\top}\boldsymbol{\Lambda}_{i}\boldsymbol{(}\boldsymbol{\theta}_{i}-\bar{\boldsymbol{\theta}}_{i})$
$=\frac{1}{2}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})^{\top}\boldsymbol{\Lambda}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})$
where $\boldsymbol{\Lambda}_{i}=\operatorname{diag}(\frac{2G_{i}I_{i}}{\ell_{i}},\frac{E_{i}I_{i}}{\ell_{i}},\frac{E_{i}I_{i}}{\ell_{i}})$, $G_{i}$, $E_{i}$, $I_{i}$ are the shear modulus, Young’s modulus, and cross-sectional inertia of the $i$th joint, respectively. The block-diagonal matrix $\boldsymbol{\Lambda}$ is composed of $\boldsymbol{\Lambda}_{i}$, referred to as the joint stiffness matrix. We also note $G_{i}=\frac{E_{i}}{2(1+\rho_{i})}$ for the isotropic material used in the MeSCR, where $\rho_{i}$ represents the Poisson’s ratio. The material properties of each joint are determined by whether it comprises a hard-magnetic material or a soft matrix, both of which exhibit a positive Poisson’s ratio.
II-B2 Magnetic Potential Energy
As shown in Fig. 2, each embedded micromagnet is treated as an independent joint, with additional joints added on either side to assist the description. The joint indices of all $N_{m}$ micromagnets constitute an index set $\mathbb{K}\subset\mathbb{Z}$. Given that the robot operates under a constant temperature, the formulation aligns with the standard Helmholtz free energy without thermal variations [10]. Additional terms with thermal effects can be found in [29]. In this case, the magnetic potential energy of the MeSCR is given by
$\mathcal{E}_{m}=-\sum_{k\in\mathbb{K}}\mathbf{m}_{k}^{\top}\mathbf{b}(\mathbf{p}_{k})$
where $\mathbf{m}_{k}=\mathbf{R}_{0}^{k}\bar{\mathbf{m}}_{k}\in\mathbb{R}^{3}$ is the magnetic dipole moments of the $k$th magnet with modulus $M_{k}=\|\bar{\mathbf{m}}_{k}\|$ and $\mathbf{b}(\mathbf{p}_{k})\in\mathbb{R}^{3}$ denote the external magnetic field with modulus $B_{k}=\|\mathbf{b}(\mathbf{p}_{k})\|$. For readibility, we sometimes denote $\mathbf{b}(\mathbf{p}_{k})$ in $\mathbf{b}_{k_{i}}$, in which $i=0,1,\dots,N_{m}-1$.
It is worth noting that some operational scenarios of MeSCR (e.g., interaction with human tissue) typically require low-speed motion to ensure operational safety. Therefore, it is generally assumed that MeSCR operates in quasi-static states, where the inertia terms can be neglected. The dynamic analysis of MeSCR is thus reduced to finding the equilibrium between internal/external forces and potential energy, which can be obtained by solving the following simplified Lagrangian equation:
$\nabla_{\boldsymbol{\theta}}\mathcal{E}_{e}+\nabla_{\boldsymbol{\theta}}\mathcal{E}_{m}+\boldsymbol{\tau}(\boldsymbol{\theta})=\mathbf{0}$ (5)
where $\boldsymbol{\tau}(\boldsymbol{\theta})$ is the internal/external applied torques that can be considered (e.g., gravity torques). We further provide the analytical expressions for the gradients and Hessians of the potential energies, which can improve the accuracy and efficiency of calculations. The gradient and Hessian of $\mathcal{E}_{e}$ are straightforward, whereas we focus on computing that of $\mathcal{E}_{m}$.
$\nabla_{\boldsymbol{\theta}_{i}}\mathcal{E}_{m}=$ $-\sum_{k\in\mathbb{K},k\geq i}\left(\left[\mathbf{m}_{k}\right]_{\times}^{\top}\mathbf{R}_{0}^{i}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}\right)^{\top}\mathbf{b}(\mathbf{p}_{k})$ (6)
$-\sum_{k\in\mathbb{K},k\geq i}\left(\frac{\partial\mathbf{b}(\mathbf{p}_{k})}{\partial\mathbf{p}_{k}}\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}\right)^{\top}\mathbf{m}_{k}$
$\triangleq$ $-\underbrace{\mathbf{M}_{i}(\boldsymbol{\theta})\mathbf{b}}_{\begin{subarray}{c}\text{magnetic}\\ \text{torques}\end{subarray}}-\sum_{k\in\mathbb{K},k\geq i}\left(\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}\right)^{\top}\underbrace{\left(\nabla_{\mathbf{p}_{k}}\mathbf{b}\right)\mathbf{m}_{k}}_{\begin{subarray}{c}\text{magnetic}\\ \text{forces}\end{subarray}}$
where $\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}$ represents the right Jacobian of $\mathrm{SO}(3)$ manifold associated with $\boldsymbol{\theta}_{i}$ [30]. We reorganize the expression by grouping the multiplicative terms into matrix $\mathbf{M}_{i}(\boldsymbol{\theta})\in\mathbb{R}^{3\times 3N_{m}}$ and introducing vector $\mathbf{b}=\left[\mathbf{b}_{k_{0}}^{\top},\dots,\mathbf{b}_{k_{N_{m}-1}}^{\top}\right]^{\top}\in\mathbb{B}\subset\mathbb{R}^{3N_{m}}$ to obtain a compact form. For precise calculations, $\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}$ can be derived from (3). However, the magnetic force is usually negligible in comparison to the magnetic torques, which implies the following explicit representation of (5) as
$\boldsymbol{\Lambda}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})+\boldsymbol{\tau}(\boldsymbol{\theta})=\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$ (7)
where $\mathbf{M}(\boldsymbol{\theta})=\left[\mathbf{M}_{0}^{\top},\dots,\mathbf{M}_{N-1}^{\top}\right]^{\top}\in\mathbb{R}^{3N\times 3N_{m}}$. By developing (7) and performing some rearrangements, the joint space compliance relation between the variation of the actuating magnetic field $(\delta\mathbf{b})$ and the configuration variation $(\delta\boldsymbol{\theta})$ is obtained:
$\underbrace{\left(\boldsymbol{\Lambda}+\frac{\partial^{2}\mathcal{E}_{m}}{\partial\boldsymbol{\theta}\partial\boldsymbol{\theta}^{\top}}+\frac{\partial\boldsymbol{\tau}}{\partial\boldsymbol{\theta}}\right)}_{\triangleq\mathbf{S}}\delta\boldsymbol{\theta}=\mathbf{M}(\boldsymbol{\theta})\delta\mathbf{b}$
where the Hessian of $\mathcal{E}_{m}$ (denoted in $\mathbf{S}_{m}$) is given in the supplementary material and $\mathbf{S}\in\mathbb{R}^{3N\times 3N}$ denotes the Hessian matrix of stiffness. Hence, the joint-space compliance relation reads
$\delta\boldsymbol{\theta}=\mathbf{S}^{-1}\mathbf{M}(\boldsymbol{\theta})\delta\mathbf{b}.$ (8)
Premultiplying (8) by the space Jacobian (2) yields the Cartesian compliance mapping from magnetic field perturbations to distal pose variations:
$\delta\mathbf{h}=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{S}^{-1}\mathbf{M}(\boldsymbol{\theta})\delta\mathbf{b}\triangleq\mathbf{J}_{\mathbf{b}}\delta\mathbf{b}$ (9)
where $\mathbf{J}_{\mathbf{b}}\in\mathbb{R}^{6\times 3N_{m}}$ is referred to as the actuation Jacobian. We close this subsection by introducing two useful propositions:
Proposition 1.
The matrix-valued function $\mathbf{M}(\boldsymbol{\theta})$ and the magnetic torque $\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$ are bounded on $\mathbb{R}^{3N}$ with positive constants $\mathcal{M}_{0}$ and $\mathcal{M}$ such that $\|\mathbf{M}(\boldsymbol{\theta})\|\leq\mathcal{M}_{0}$ and $\|\mathbf{M}(\boldsymbol{\theta})\mathbf{b}\|\leq\mathcal{M}$ for all $\boldsymbol{\theta}\in\mathbb{R}^{3N}$, respectively.
Proposition 2.
The vector-valued function $\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$ is Lipschitz continuous on $\mathbb{R}^{3N}$ with a Lipschitz constant $\mathcal{L}$ such that $\|\mathbf{M}(\boldsymbol{\theta})\mathbf{b}-\mathbf{M}(\boldsymbol{\varphi})\mathbf{b}\|\leq\mathcal{L}\|\boldsymbol{\theta}-\boldsymbol{\varphi}\|$ for all $\boldsymbol{\theta},\boldsymbol{\varphi}\in\mathbb{R}^{3N}$.
Proof.
The proofs are deferred to Appendix A. ∎
II-C Analysis and Validation of Equilibrium Solution
[…]
TABLE I: Geometric and Material properties of the MeSCR#1
[…]
$I$ (Area moment of inertia) $0.7854$ $[\mbox{mm}^{4}]$
[…]
$M$ (Modulus of the magnetic dipole moment) $10^{-2}$ $[\mbox{A}\cdot\mbox{m}^{2}]$
We first analyze the equilibrium solution. Assuming the robot to be naturally straight $(\bar{\boldsymbol{\theta}}=\mathbf{0})$, and $\boldsymbol{\tau}(\boldsymbol{\theta})$ is ruled out. Then, reformulate (7) derives the following fixed-point equation:
$\boldsymbol{\theta}=\boldsymbol{\Lambda}^{-1}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}.$ (10)
Observe that the right-hand side of the equation represents a continuous mapping with respect to $\boldsymbol{\theta}$. We present the following theorem regarding the uniqueness of the solution.
Theorem 1.
There exists a unique equilibrium solution to (10) if the following inequality holds:
$\max_{k\in\mathbb{K}}B_{k}<\frac{\pi}{4}\frac{\lambda_{\min}\left(\boldsymbol{\Lambda}\right)}{\sum_{k\in\mathbb{K}}kM_{k}}$ (11)
where $\lambda_{\min}(\cdot)$ denotes the smallest eigenvalue of a matrix.
Proof.
Based on the Proposition 2, the Lipschitz constant of the continuous mapping can be estimated by
$\left\|\boldsymbol{\Lambda}^{-1}\frac{\partial}{\partial\boldsymbol{\theta}}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}\right\|=\|\boldsymbol{\Lambda}^{-1}\mathbf{S}_{m}\|\leq\|\boldsymbol{\Lambda}^{-1}\|\mathcal{L}=\lambda_{\min}^{-1}(\boldsymbol{\Lambda})\mathcal{L}.$
From (11), it follows that $\lambda_{\min}^{-1}(\boldsymbol{\Lambda})\mathcal{L}<1$. Then, $\boldsymbol{\Lambda}^{-1}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$ is a contraction mapping on $\mathbb{R}^{3N}$. By Banach’s fixed-point theorem, there exists a unique equilibrium point to (10). ∎
The inequality provides a sufficient condition to guarantee the uniqueness of a globally attractive equilibrium point for any initial condition. Notably, the upper bound is solely determined by the material properties: $E$, $I$, $L$, $M_{k}$, and is independent of the joint number $N$ (note that $N\bar{\ell}_{i}\approx L)$. The limitation on the external magnetic field strength can be explained by the critical buckling instability of the magneto-elastic material [31]. We found it intriguing that when each joint is magnetized, the resulting upper bound exhibits a striking structural resemblance to that derived in our previous work [32] on magnetic continua, differing only slightly in the coefficients. This similarity highlights the underlying consistency in the physical principles governing continuum and discreteness. Furthermore, Theorem 1 also ensures $\mathbf{S}$ to be invertible. By applying Weyl’s inequality, one can easily verify that $\lambda_{\min}(\mathbf{S})\geq\lambda_{\min}(\boldsymbol{\Lambda})-\mathcal{L}>0$.
Although the upper bound remains a conservative estimate, once the external magnetic field strength exceeds this upper limit, (10) may admit multiple equilibria pertaining to different initial conditions [21]. Therefore, selecting appropriate initial values for the iteration becomes crucial to ensuring convergence and avoiding undesirable behaviors. The solution trajectory can be described by the constructed dynamic system $\frac{\mathrm{d}\boldsymbol{\theta}}{\mathrm{d}t}=-\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}}$ with initial condition $\boldsymbol{\theta}(0)$, which can also be regarded as a gradient flow that minimizes the energy function $\mathcal{E}(\boldsymbol{\theta})$. We propose the following theorem and lemmas, which guarantee that the solution is bounded and convergent by choosing an appropriate initial condition.
Lemma 1.
For any initial condition $\boldsymbol{\theta}(0)$, the solution trajectory $\boldsymbol{\theta}(t)$ is bounded that
$\|\boldsymbol{\theta}(t)\|\leq\max\left\{\|\boldsymbol{\theta}(0)\|,\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\right\}.$
Proof.
Consider the Lyapunov function $V(\boldsymbol{\theta})=\frac{1}{2}\|\boldsymbol{\theta}\|^{2}$, the derivative of $V(\boldsymbol{\theta})$ with respect to time is $\dot{V}(\boldsymbol{\theta})=\boldsymbol{\theta}^{\top}(-\boldsymbol{\Lambda}\boldsymbol{\theta}+\mathbf{M}(\boldsymbol{\theta})\mathbf{b})=-\boldsymbol{\theta}^{\top}\boldsymbol{\Lambda}\boldsymbol{\theta}+\boldsymbol{\theta}^{\top}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$. Using Proposition 1, the upper bound of $\dot{V}(\boldsymbol{\theta})$ is derived as $\dot{V}(\boldsymbol{\theta})\leq-\lambda_{\min}(\boldsymbol{\Lambda})\|\boldsymbol{\theta}\|^{2}+\mathcal{M}\|\boldsymbol{\theta}\|.$ Let $W(t)=\sqrt{2V(\boldsymbol{\theta})}=\|\boldsymbol{\theta}\|$, one can find out that: $\dot{W}(t)=\frac{\dot{V}(\boldsymbol{\theta})}{W(t)}\leq-\lambda_{\min}W(t)+\mathcal{M}.$ Applying Grönwall’s inequality, we obtain $W(t)\leq W(0)e^{-\lambda_{\min}(\boldsymbol{\Lambda})t}+\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\left(1-e^{-\lambda_{\min}\left(\boldsymbol{\Lambda}\right)t}\right).$ Observed that as $t\to\infty$, the exponential term vanishes, leading to $\lim_{t\to\infty}W(t)\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}.$ This completes the proof, establishing that $\|\boldsymbol{\theta}(t)\|$ remains $\|\boldsymbol{\theta}(t)\|\leq\max\left\{\|\boldsymbol{\theta}(0)\|,\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\right\}.$
Lemma 2.
Let $\boldsymbol{\theta}^{*}$ be any equilibrium point of (10). Using Proposition 1, we have $\|\boldsymbol{\theta}^{*}\|=\|\boldsymbol{\Lambda}^{-1}\mathbf{M}(\boldsymbol{\theta}^{*})\mathbf{b}\|\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}$.
Theorem 2.
The solution trajectory $(\boldsymbol{\theta}(t))$ will always converge to the largest invariant subset of the equilibrium point set if the initial condition satisfies $\|\boldsymbol{\theta}(0)\|\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}$.
Proof.
Choose the Lyapunov function $V(\boldsymbol{\theta})=\mathcal{E}(\boldsymbol{\theta})$, whose derivative along the solution trajectory satisfies $\dot{V}(\boldsymbol{\theta})=-\|\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}}\|^{2}\leq 0$, with equality holding if and only if $\boldsymbol{\theta}$ is an equilibrium point. Consider the positively invariant set $\Omega=\left\{\boldsymbol{\theta}\in\mathbb{R}^{3N}\mid\|\boldsymbol{\theta}\|\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\right\}$ (Lemma 1), which contains all equilibrium points (Lemma 2). By Brouwer’s fixed-point theorem, there exists at least one equilibrium point within $\Omega$, that is, $\dot{V}(\boldsymbol{\theta})\leq 0$ holds for all $\boldsymbol{\theta}\in\Omega$. According to LaSalle’s invariance principle, every trajectory within $\Omega$ converges to the largest invariant subset where $\dot{V}(\boldsymbol{\theta})=0$. ∎
When the external magnetic field changes, requiring the recalculation of the equilibrium point, Theorem 2 provides a natural choice for the initial condition as $\boldsymbol{\theta}(0)=\mathbf{0}$ because $\forall\mathbf{b}\in\mathbb{B},\mathcal{M}\geq 0$. One can also choose the nearby solution as an initial guess if $\mathbf{b}$ continuously increases. For small perturbations on $\mathbf{b}$, the corresponding $\delta\boldsymbol{\theta}$ can be calculated by (8). Moreover, from Proposition 1 and the fact that $\boldsymbol{\theta}=\mathbf{0}$ is the unique equilibrium solution when $\mathbf{b}=\mathbf{0}$, we deduce that $\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}}$ is Lipschitz continuous with respect to $\mathbf{b}$, and the equilibrium points depends continuously on it.
[…]
Notably, using fewer joints does not introduce significant truncation errors and offers a computational speed advantage over BVP-solving models. Fig. 3e and Fig. 3f show the root-mean-square error (RMSE) of the centerline position and the distal pose error norm as the joint count increases from $3$ to $100$, benchmarked against $200$ joints. Results demonstrate that RMSE decreases proportionally to $1/N^{2}$, with an error bound established in our previous work [34]. Using $7$ joints achieves an RMSE below the acceptable $10^{-4}$ threshold. Additionally, the distal pose error norm declines rapidly as $N$ grows, confirming that fewer joints can accurately compute the distal pose under external magnetic fields. This analysis leads to a key insight: the performance index of the actuation Jacobian is largely insensitive to $N$ and can be understood as an inherent property of the robot in a specific configuration.
III Enhancing Kinematic Performances
This section focuses on the structural optimization of MeSCRs to enhance kinematic performances, including manipulability and dexterity. Prior to the optimization, we perform some kinematic analysis of the actuation Jacobian, uncovering how different magnet configuration shape the robot’s controllable DoF achievable workspace.
III-A Kinematics Analysis
We define the controllable degrees of freedom (DoF) as the independently actuated direction of the task-space twist, which corresponds to the row rank of the actuation Jacobian ($\mathbf{J}_{\mathbf{b}}$). The following theorem reveals the relation between it and the configuration of embedded magnets:
Theorem 3.
The controllable DoF of the MeSCR does not exceed twice the number of embedded magnets.
Proof.
Recall that the actuation Jacobian $\mathbf{J}_{\mathbf{b}}=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{S}^{-1}\mathbf{M}(\boldsymbol{\theta})$, the post-multiplied matrix $(\mathbf{M}(\boldsymbol{\theta}))$ can be rewritten as
$\mathbf{M}(\boldsymbol{\theta})$ $=\mathbf{D}_{k_{N_{m}-1}}^{\top}\left(\begin{bmatrix}\begin{matrix}\mathbf{1}_{k_{0}}\\ \mathbf{0}_{N-k_{0}}\end{matrix}\cdots\begin{matrix}\mathbf{1}_{k_{N_{m}-1}}\\ \mathbf{0}_{N-k_{N_{m}-1}}\end{matrix}\end{bmatrix}\otimes\mathbf{I}_{3\times 3}\right)$ (12)
×$\times\operatorname{blkdiag}\left\{\left[\mathbf{m}_{k_{0}}\right]_{\times},\dots,\left[\mathbf{m}_{k_{N_{m}-1}}\right]_{\times}\right\}$
in which the block diagonal matrix $\mathbf{D}_{k_{N_{m}-1}}$ is provided in (24) and $\mathbf{1}_{k}$ denotes the all-ones vector of $k$ dimensions. To determine the rank of $\mathbf{M}(\boldsymbol{\theta})$, one starts by noting that
$\operatorname{rk}\left(\begin{bmatrix}\begin{matrix}\mathbf{1}_{k_{0}}\\ \mathbf{0}_{N-k_{0}}\end{matrix}\cdots\begin{matrix}\mathbf{1}_{k_{N_{m}-1}}\\ \mathbf{0}_{N-k_{N_{m}-1}}\end{matrix}\end{bmatrix}\otimes\mathbf{I}_{3}\right)=3N_{m}.$
Applying Sylvester’s rk inequality, we obtain:
$\operatorname{rk}\left(\mathbf{M}(\boldsymbol{\theta})\right)$ $\leq\min\!\left\{\sum_{i=0}^{N_{m}-1}\!\operatorname{rk}\!\left(\mathbf{R}_{0}^{k_{i}}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k_{i}}}\right),\sum_{i=0}^{N_{m}-1}\!\operatorname{rk}\!\left(\left[\mathbf{m}_{k_{i}}\right]_{\times}\right)\right\}$ (13)
$=2N_{m},$
and
$\operatorname{rk}\!\left(\mathbf{M}(\boldsymbol{\theta})\right)$ $\geq\sum_{i=0}^{N_{m}-1}\!\left(\operatorname{rk}\!\left(\mathbf{R}_{0}^{k_{i}}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k_{i}}}\right)+\!\operatorname{rk}\!\left(\left[\mathbf{m}_{k_{i}}\right]_{\times}\right)\right)-3N_{m}$ (14)
$=2N_{m}.$
Here we used the facts that $\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k_{i}}}$ and $\mathbf{R}_{0}^{k_{i}}$ are both invertible and the rank of a skew-symmetric matrix remains $2$. Combined (13) with (14) leads to the following result:
$\operatorname{rk}\left(\mathbf{M}(\boldsymbol{\theta})\right)=2N_{m}.$
Consequently, an important conclusion is derived as $\operatorname{rk}\left(\mathbf{J}_{\mathbf{b}}\right)\leq\min\left\{6,\operatorname{rk}\left(\mathbf{M}(\boldsymbol{\theta})\right)\right\}=\min\left\{6,2N_{m}\right\}$. ∎
We proceed with a brief discussion on MeSCR’s structural configuration with different $N_{m}$:
Figure 4: Deformation of the MeSCR with different numbers of embedded magnets $(N_{m}=1,2,3,4)$, corresponding to subfigures a.), b.), c.), and d.). As $N_{m}$ increases, the MeSCR is characterized as underactuated for $N_{m}=1,2$, fully-actuated for $N_{m}=3$, and redundant for $N_{m}=4$.
  • $N_{m}=1$ [2, 3, 4] (see Fig. 4a.): The simplest yet most used configuration. The single magnet is typically embedded at the distal end to maximize torque transmission. From Theorem 3, we show that at most $2$ DoF of the task-space twist is controllable, which is consistent with the conclusion in [18].
  • $N_{m}=2$[35, 20, 19, 36] (see Fig. 4b.): Multimodality properties emerged since the orientation of embedded magnets is designable. However, at most $4$ DoF are controllable, which still characterizes the system as underactuated.
  • $N_{m}=3$ [37, 13, 38] (see Fig. 4c.): The structural configuration of the embedded magnets is also optimizable. The maximum controllable DoF is $6$, indicating that the MeSCR is fully-actuated.
  • $N_{m}\geq 4$[39, 40, 41] (see Fig. 4d.): Adopted by many commerical products. The column rank of $\mathbf{J}_{\mathbf{b}}$ exceeds its row rank, indicating that the MeSCR is a redundant system and the number of redundant DoF is $62N_{m}-$ 6.
The maximum controllable DoF will also be limited by the actuating field. As $N_{m}$ increases, achieving fully-actuated or redundant control requires each embedded magnet to be paired with an independent actuating field. However, generating independently controllable actuating fields at arbitrary positions within the workspace is highly challenging. While some advanced equipment may be capable of this [42], it is not feasible for most conventional systems. For example, when the actuating field is spatially uniform within the workspace, which can be expressed as $\mathbf{b}=\mathbf{U}_{u}\mathbf{b}_{u}$ with $\mathbf{U}_{u}=\mathbf{1}_{N_{m}}\otimes\mathbf{I}_{3}$ and $\mathbf{b}_{u}\in\mathbb{R}^{3}$. In this case, the maximum controllable DoF will not exceed $3$. To reduce the number of design variables, we assume all embedded magnets have uniform material and geometric properties, which also facilitates the manufacturing process. Consequently, the orientation and position of the embedded magnets become the primary variables to be optimized.
In existing commercial products, embedded magnets are generally placed along the catheter’s axial direction. For $N_{m}$ magnets, this arrangement leads to $2^{N_{m}}$ possible permutations. However, if the actuating magnetic field of each magnet is independently controllable, the directions of the magnetic moments become redundant. This redundancy arises because of the relationship: [±$[\pm\mathbf{m}_{k}]_{\times}\mathbf{b}_{k}=[\mathbf{m}_{k}]_{\times}(\pm\mathbf{b}_{k})$, indicating that the effect of flipping a magnetic moment’s direction can always be compensated by reversing the corresponding magnetic field direction. Based on this reasoning, we assume that the magnetic field can only undergo an overall direction reversal. For systems driven by electromagnets, this can be easily achieved by altering the current direction.
Aligning the magnetic moments along the axial direction significantly reduces the optimization search space but imposes a limitation on the motion of the MeSCR. Specifically, this arrangement restricts the MeSCR to material-twist-free due to the decoupling of twisting and bending, as ensured by the isotropic and axially symmetric properties of its soft matrix.
Theorem 4 (Material-twist-free).
The MeSCR is material-twist-free ($\boldsymbol{\theta}_{i}^{\top}\bar{\mathbf{t}}_{i}=0$) for any actuating magnetic field, provided that the embedded magnets’ magnetic moments are aligned axially, i.e., $\bar{\mathbf{m}}_{k}=\pm M_{k}\bar{\mathbf{t}}$, and the magnetic field strength satisfies:
$\max_{k\in\mathbb{K}}B_{k}\leq\frac{\sqrt{6}\lambda_{\min}(\boldsymbol{\Lambda})}{\sum_{k\in\mathbb{K}}M_{k}}.$ (15)
Proof.
The proof is detailed in the supplementary material. ∎
Noticing that $\frac{\sqrt{6}\lambda_{\min}(\boldsymbol{\Lambda})}{\sum_{k\in\mathbb{K}}M_{k}}\propto N^{2}$, as $N\to\infty$, Theorem 4 holds for any actuating magnetic field. This upper bound characterizes the ability of the discrete model with $N$ joints to approximate the continuum, as discussed in Section II-C. We found an alternative form of the equilibrium equation (10) from the underlying proof:
$\boldsymbol{\theta}_{i}^{*}=\frac{\sum_{k\geq i}\mathbf{v}_{k}}{\lambda_{\max}(\boldsymbol{\Lambda}_{i})}=\frac{\sum_{k\geq i}[\mathbf{R}_{0}^{k^{\top}}(\boldsymbol{\theta}^{*})\mathbf{b}_{k}]_{\times}^{\top}\bar{\mathbf{m}}_{k}}{\lambda_{\max}(\boldsymbol{\Lambda}_{i})}.$ (16)
For any initial condition satisfying the assumptions of Theorem 2, the associated equilibrium state is guaranteed to exhibit zero torsional components. The configuration space (1) thus can be degenerated to a simplified model as presented in [16]. Moreover, the joint stiffness matrix reduces to $\boldsymbol{\Lambda}_{i}=\Lambda_{i}\mathbf{I}_{3}$ with $\Lambda_{i}=\frac{E_{i}I_{i}}{\ell_{i}}$, and we denote $\lambda_{\min}(\boldsymbol{\Lambda})=\Lambda_{\min}$. We emphasize its distinction from geometric-twist-free, which requires $\boldsymbol{\theta}_{i}^{\top}\mathbf{t}_{i}=0$. It should be known that geometric torsion may still exist under the material-twist-free condition. However, if the external magnetic field is coplanar, the following corollary holds:
Corollary 1.
Define the plane $\Pi=\operatorname{span}\{\bar{\mathbf{t}},\mathbf{v}\}\subset\mathbb{R}^{3}$, where $\mathbf{v}\in\mathbb{R}^{3}$ is linearly independent of $\bar{\mathbf{t}}$. Suppose $\forall k\in\mathbb{K}$, $\bar{\mathbf{m}}_{k}=\pm M_{k}\bar{\mathbf{t}}$ and $\mathbf{b}_{k}\in\Pi$. Then there exists an equilibrium configuration $\boldsymbol{\theta}^{*}\in\mathbb{R}^{3N}$ such that the centerline of the MeSCR lies entirely within the plane $\Pi$.
Proof.
[…]
Specially, if the initial value is chosen as $\boldsymbol{\theta}(0)=\mathbf{0}$, then the gradient-based iteration evolves within the subspace $\operatorname{span}\{\mathbf{n}\}$. By Theorem 2, the optimization trajectory will converge to a planar equilibrium configuration satisfying $\boldsymbol{\theta}_{i}=\theta_{i}\mathbf{n}$. We can further derive a closed-form expression under the small-angle premise, wherein the cumulative rotation matrix admits the first-order expansion: $\mathbf{R}\left(\sum_{i=0}^{k}\theta_{i}\mathbf{n}\right)\approx\mathbf{I}_{3}+\sum_{i=0}^{k}\theta_{i}[\mathbf{n}]_{\times}$. Substituting this into the total potential energy yields
$\mathcal{E}\approx\sum_{i=0}^{N-1}\frac{1}{2}\Lambda_{i}\theta_{i}^{2}+\sum_{i=0}^{N-1}\theta_{i}\!\left(\sum_{k\in\mathbb{K},k\geq i}\bar{\mathbf{m}}_{k}^{\top}[\mathbf{n}]_{\times}\mathbf{b}_{k}\right)-\sum_{k}\bar{\mathbf{m}}_{k}^{\top}\mathbf{b}_{k}.$
The optimization problem $\min_{\boldsymbol{\theta}}\mathcal{E}$ thus has the explicit solution
$\theta_{i}=\frac{1}{\Lambda_{i}}\sum_{k\in\mathbb{K},k\geq i}\bar{\mathbf{m}}_{k}^{\top}[\mathbf{b}_{k}]_{\times}\mathbf{n}.$ (17)
It can be verified that this also corresponds to the general solution of (16) under the small-angle premise.
III-B Optimization Methodology
Figure 5: Comparison of the half-planar normalized workspace of MeSCRs with equidistant placement of three embedded magnets.
The position and orientation of embedded magnets determine the magnetic response deformation of MeSCRs under the same magnetic field. For instance, embedded magnets with variable positions have been employed to expand the MeSCR’s workspace [8]. In the PRB model, the length of embedded magnets can be defined by the Voronoi length as $L_{k}\triangleq\sum_{i\leq k}\ell_{i},\forall k\in\mathbb{K}$. From (10), it follows that the flexible segment beyond the last embedded magnet is unactuated. Consequently, it does not affect the workspace volume of the MeSCR, nor any performance metric defined using a left-invariant Riemannian metric. Therefore, one of the magnets is fixed at the distal end of the continuum ($L_{k_{N_{m}-1}}=L$), leaving $N_{m}-1$ decision variables. Additionally, let the flexible segments between adjacent magnets be uniformly distributed and have appropriate spacing to mitigate magnetic coupling effects. Based on the previous assumption regarding the overall reversal of the magnetic field direction, any configuration of axially placed magnets has a corresponding symmetrical configuration, resulting in a total of $2^{N_{m}-1}$ possible combinations.
Expanding the workspace volume is a key objective in the optimal design of MeSCRs [43]. However, comparing workspace volumes across MeSCRs with different embedded magnet orientations is of limited significance, unless the focus is explicitly on spatial coverage by the distal end. As shown in Fig. 5, we compared the normalized half-workspace areas of the four MeSCRs under equidistant configurations. The results indicate that ② achieves the largest normalized workspace area, whereas ③ has the smallest. Nevertheless, despite its smaller workspace area, ① may still reach regions that are inaccessible to ②. Therefore, our focus shifts to the global indices of MeSCRs with fixed magnet orientations within their respective workspaces.
Recall that ${}_{s}^{b}\mathbf{H}(\boldsymbol{\theta}):\mathbb{T}\to\mathrm{SE}(3)$ is the forward kinematic map of the MeSCR. Let $z({}_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k}):\mathbb{T}\to\mathbb{R}$ be an integrable scalar functional defined on $\mathbb{T}$, representing a local performance density of ${}_{s}^{b}\mathbf{H}(\boldsymbol{\theta})$ with left-invariant Riemannian metric. This local measure can be extended to a global measure $\mathcal{Z}(L_{k})$ by integrating $z$ over $\mathbb{T}$, yielding the following optimization problem on $L_{k}$:
$\min_{L_{k}}$ $\ \mathcal{Z}(L_{k})\triangleq\int_{\mathbb{T}}z({}_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\ \Omega_{\mathbb{T}}$ (18)
[…]
where $\Omega_{\mathbb{T}}$ represents the Riemannian volume form on $\mathbb{T}$ induced from the Euclidean metric and $\mathcal{C}(L_{k})$ denotes linear constraints on $L_{k}$. The equilibrium constraint restricts feasible configurations to the solution set $\mathbb{T}^{*}\triangleq\{\boldsymbol{\theta}\in\mathbb{T}|\nabla_{\boldsymbol{\theta}}\mathcal{E}(\boldsymbol{\theta},\mathbf{b})=\mathbf{0}\}$. If the actuation manifold $\mathbb{B}$ is arbitrary, the set $\mathbb{T}^{*}$ may not admit an explicit parametrization. Nevertheless, under mild regularity conditions on $\mathcal{E}$, and by Sard’s theorem, the solution set $\mathbb{T}^{*}\subset\mathbb{T}$ is countably $d$-rectifiable, with $d=3N-\operatorname{rk}(\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}})$. Hence, it admits a well-defined $\mathcal{H}^{d}$-measure and supports geometric integration. Consequently, the constrained integral over $\mathbb{T}$ may be rewritten as an iterated integral over $\mathbb{B}$ and the corresponding solution manifold:
$\mathcal{Z}(L_{k})=\int_{\mathbb{B}}\left(\int_{\mathbb{T}^{*}}z(_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\mathrm{d}\mathcal{H}^{d}(\boldsymbol{\theta})\right)\Omega_{\mathbb{B}}.$
Now, suppose the external magnetic field $\mathbf{b}$ is spatially uniform with $\mathbf{b}=\mathbf{U}_{u}\mathbf{b}_{u}$ and satisfies the condition in Theorem 1. The manifold $\mathbb{B}$ is now an affine subspace diffeomorphic to $\mathbb{R}^{3}$. By Theorem 1, the stiffness matrix $\mathbf{S}$ is nonsingular in a neighborhood of each $\mathbf{b}_{u}\in\mathbb{B}$. By Theorem 3, the Jacobian $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}}=\mathbf{S}^{-1}\mathbf{M}$ satisfies $\operatorname{rk}(\mathbf{S}^{-1}\mathbf{M})=2N_{m}$, which ensures that $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}=\mathbf{S}^{-1}\mathbf{M}\mathbf{U_{u}}$ is of full rank provided $N_{m}\geq 2$. In this regime, the equilibrium constraint gives rise to a smooth immersion $\boldsymbol{\theta}^{*}:\mathbb{B}\to\mathbb{T}$. The inner integral reduces to an evaluation at $\boldsymbol{\theta}^{*}(\mathbf{b})$, and using the area formula [44] yields:
$\int_{\mathbb{T}^{*}}z(_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\mathrm{d}\mathcal{H}^{d}(\boldsymbol{\theta})=z(_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\mathcal{J}_{\boldsymbol{\theta}}(\mathbf{b}),$
where $\mathcal{J}_{\boldsymbol{\theta}}(\mathbf{b})$ denotes the Jacobian of the immersion $\boldsymbol{\theta}^{*}(\mathbf{b})$ and is given by
$\mathcal{J}_{\boldsymbol{\theta}}=\left[\det\left(\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}^{\top}\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}\right)\ \right]^{\frac{1}{2}}\triangleq\left[\mbox{det}\left(\mathbf{G}_{\boldsymbol{\theta}}\right)\right]^{\frac{1}{2}}.$
where $\mathbf{G}_{\boldsymbol{\theta}}$ denotes the Gram matrix of $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}$. Substituting back, the global optimization problem (18) takes the following unconstrained reformulation:
$\min_{L_{k}}$ $\ \mathcal{Z}(L_{k})=\int_{\mathbb{B}}z({}_{s}^{b}\mathbf{H}(\boldsymbol{\theta}^{*}(\mathbf{b}));L_{k})\mathcal{J}_{\boldsymbol{\theta}}\ \Omega_{\mathbb{B}}.$ (19)
The optimality is ensured by satisfying the following first-order necessary condition:
$\nabla_{L_{k}}\mathcal{Z}$ $=\int_{\mathbb{B}}\left(z\nabla_{L_{k}}\mathcal{J}_{\boldsymbol{\theta}}+\mathcal{J}_{\boldsymbol{\theta}}\nabla_{L_{k}}z\right)\ \Omega_{\mathbb{B}}$
$=\int_{\mathbb{B}}\left[\frac{1}{2}z\mathcal{J}_{\boldsymbol{\theta}}\operatorname{tr}\left(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{L_{k}}\mathbf{G}_{\boldsymbol{\theta}}\right)+\mathcal{J}_{\boldsymbol{\theta}}\nabla_{L_{k}}z\right]\ \Omega_{\mathbb{B}}$
$=0,\ \forall k\in\mathbb{K}/\{N_{m}-1\}.$
Under regularity conditions, this implies the strong-form pointwise optimality condition:
$\left(\frac{1}{2}z\operatorname{tr}\left(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{L_{k}}\mathbf{G}_{\boldsymbol{\theta}}\right)+\nabla_{L_{k}}z\right)\mathcal{J}_{\boldsymbol{\theta}}=0,\ \forall\mathbf{b}_{u}\in\mathbb{B}.$ (20)
Observed that the condition trivially holds when $\mathcal{J}_{\boldsymbol{\theta}}=0$. However, the collapse of immersion Jacobian results in a degenerate configuration. That is, the embedded magnets may have merged or be unable to generate torques. Since $\mathcal{J}_{\boldsymbol{\theta}}$ contains complex Hessian terms, we first derive a tractable approximation to facilitate further analysis.
Lemma 3.
Suppose the inequality condition in Theorem 1 holds. Then the immersion Jacobian admits the expansion
$\mathcal{J}_{\boldsymbol{\theta}}^{2}$ $=\det\left(\mathbf{U}_{u}^{\top}\mathbf{M}^{\top}\boldsymbol{\Lambda}^{-2}\mathbf{M}\mathbf{U}_{u}\right)+\epsilon$ (21)
$\triangleq\tilde{\mathcal{J}}_{\boldsymbol{\theta}}^{2}+\epsilon$
where the remainder term $\epsilon$ is bounded above as
$|\epsilon|\leq\frac{3(2\mathcal{L}-\mathcal{L}^{2})}{(1-\mathcal{L})^{2}}\frac{N_{m}^{3}\mathcal{M}_{0}^{6}}{\Lambda_{\min}^{6}},$
and satisfies $\lim_{B_{k}\to 0}|\epsilon|\to 0$.
Proof.
The proof is deferred to the supplementary material. ∎
It shows that $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$ provides a close approximation of $\mathcal{J}_{\boldsymbol{\theta}}$ in the weak-field regime, with the residual term converging to zero as the magnetic field strength decreases. We can now turn to analyze the approximation of the objective $\tilde{\mathcal{Z}}(L_{k})\triangleq\int_{\mathbb{B}}z\ \tilde{\mathcal{J}}_{\boldsymbol{\theta}}\ \Omega_{\mathbb{B}}$. Let $\mathbf{M}(\boldsymbol{\theta})$ in (12) be represented as $\mathbf{M}=\mathbf{D}_{k_{m}-1}^{\top}\mathbf{U}_{m}\mathbf{D}_{m}$, we notice
$\mathbf{U}_{m}^{\top}\mathbf{D}_{k_{N_{m}-1}}\boldsymbol{\Lambda}^{-2}\mathbf{D}_{k_{N_{m}-1}}^{\top}\mathbf{U}_{m}$ $=\mathbf{U}_{m}^{\top}\boldsymbol{\Lambda}^{-2}\mathbf{U}_{m}=\boldsymbol{\Lambda}_{m}\!\otimes\!\mathbf{I}_{3}$
where
$\boldsymbol{\Lambda}_{m}=\begin{bmatrix}\Lambda^{m}_{k_{0}},&\Lambda^{m}_{k_{0}},&\dots&\Lambda^{m}_{k_{0}}\\ \Lambda^{m}_{k_{0}},&\Lambda^{m}_{k_{1}},&\dots&\Lambda^{m}_{k_{1}}\\ \vdots&\vdots&\ddots&\vdots\\ \Lambda^{m}_{k_{0}},&\Lambda^{m}_{k_{1}},&\dots&\Lambda^{m}_{k_{N_{m}-1}}\end{bmatrix}\in\mathbb{R}^{N_{m}\times N_{m}}$
with $\Lambda_{k_{i}}^{m}=\sum_{j=0}^{k_{i}}\frac{\ell_{j}^{2}}{E_{j}^{2}I_{j}^{2}}\approx\frac{1}{k_{0}E^{2}I^{2}}\sum_{j=0}^{i}(L_{k_{j}}-L_{k_{j-1}})^{2}$ as a function of $L_{k}$. The following lemma shows the relation between the singularity of $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$ and the configuration of $L_{k}$:
Lemma 4.
Suppose the embedded magnets are placed at the two boundaries (proximal and distal ends), that is $\forall 0\leq N_{k}<N_{m}-1$, $L_{k_{N_{k}}}=0$ and $L_{k_{N_{k}+1}}=L$. Then $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}=0$.
Proof.
[…]
$\left(\mathbf{v}^{\top}\otimes\mathbf{I}_{3}\right)\mathbf{U}_{u}\mathbf{D}_{m}=\sum_{i=N_{k}+1}^{N_{m}-1}\left[\mathbf{m}_{k_{i}}\right]_{\times}=\left[\sum_{i=N_{k}+1}^{N_{m}-1}\mathbf{m}_{k_{i}}\right]_{\times}.$
Consequently, the rank of the expression for $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$ shows that
$\operatorname{rk}\!\left(\mathbf{U}_{u}^{\top}\mathbf{D}_{m}^{\top}\!\left(\boldsymbol{\Lambda}_{m}\!\otimes\mathbf{I}_{3}\right)\mathbf{U}_{u}\mathbf{D}_{m}\right)\!$
$=\operatorname{rk}\!\left(\!\left(\mathbf{v}^{\top}\!\!\otimes\mathbf{I}_{3}\right)\!\mathbf{U}_{u}\mathbf{D}_{m}\right)=\operatorname{rk}\left(\left[\sum_{i=N_{k}+1}^{N_{m}-1}\mathbf{m}_{k_{i}}\right]_{\times}\right)\leq 2.$
That is, the actuation mapping becomes deficient in rank, and we conclude that $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}=0$. ∎
Lemma 4 indicates that placing all embedded magnets at boundary positions leads to a singularity in $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$, as expected. This degeneration of the measure can also be interpreted in terms of an effective reduction in the number of independent magnets. Specifically, magnets located at the proximal end have no actuation effect, while those at the distal end act collectively as a single unit. As a result, under boundary-concentrated configurations, the effective number of embedded magnets is reduced to one. From Theorem 3, we notice $\operatorname{rk}\left(\mathbf{M}(\boldsymbol{\theta})\right)=2$ when $N_{m}=1$, and the Jacobian $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}$ becomes deficient in rank, leading to a singularity of immersion Jacobian. This structural degeneracy provides a tractable mechanism for determining the optimal solution to a class of objective functions, as formalized in the following theorem.
Theorem 5.
If the performance index satisfies $z(\mathbf{b}_{u};L_{k})>0$, and the structural parameter $L_{k}$ satisfies the boundary configuration condition in Lemma 4, then the objective $\tilde{\mathcal{Z}}(L_{k})$ attains its minimum at this configuration.
Proof.
Since $z>0$ and $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}\geq 0$, it follows that $\tilde{\mathcal{Z}}(L_{k})\geq 0$ for any admissible $L_{k}$. When $L_{k}$ satisfies the boundary configuration condition in Lemma 4, we have $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}(\mathbf{b})\equiv 0$ over $\mathbb{B}$, and thus $\tilde{\mathcal{Z}}(L_{k})=0$. ∎
Due to the Jacobian singularity at boundary configurations, the value of the objective becomes insensitive to the gradient of the performance index, thereby weakening the meaningfulness of the solution. However, the minimum spacing constraint prevents the structural parameter $L_{k}$ from reaching such degenerate configurations. Since $\mathcal{J}_{\boldsymbol{\theta}}$ remains continuous on $L_{k}$, it stays small in the vicinity of the boundary, making nearby configurations highly competitive in minimizing the objective. In this case, the performance gradient becomes the key factor in selecting specific configurations. One can also reduce the optimization bias by considering the objective as $\mathcal{Z}/\mathcal{V}$, where $\mathcal{V}\triangleq\int_{\mathbb{B}}\mathcal{J}_{\boldsymbol{\theta}}\Omega_{\mathbb{B}}$ is the total induced volume. This normalized objective reflects the average performance on the configuration space and is insensitive to Jacobian singularity.
Now we focus on the nontrivial solution of (20). If $z$ is a non-zero performance index, one can further deduce that
$\operatorname{tr}(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{L_{k}}\mathbf{G}_{\boldsymbol{\theta}})=\sum_{i}\frac{\nabla_{L_{k}}\lambda_{i}(\mathbf{G}_{\boldsymbol{\theta}})}{\lambda_{i}(\mathbf{G}_{\boldsymbol{\theta}})}=-2\frac{\nabla_{L_{k}}z}{z}.$ (22)
Geometrically, this condition implies that optimizing the local performance requires structurally modulating the spectrum of the configuration space metric to counteract its distortion. From Lemma 3, we have the following approximation:
$\mathbf{G}_{\boldsymbol{\theta}}$ $\approx\mathbf{U}_{u}^{\top}\mathbf{D}_{m}^{\top}\left(\boldsymbol{\Lambda}_{m}\otimes\mathbf{I}_{3}\right)\mathbf{D}_{m}\mathbf{U}_{u}$ (23)
$=\sum_{ij}\Lambda_{ij}^{m}\left(\mathbf{m}_{k_{i}}\mathbf{m}_{k_{j}}^{\top}-\left(\mathbf{m}_{k_{i}}^{\top}\mathbf{m}_{k_{j}}\right)\mathbf{I}_{3}\right)$
in which $\mathbf{m}_{k}=M(\cos(\theta_{0}^{k})\bar{\mathbf{t}}-\sin(\theta_{0}^{k})\mathbf{v})$. While the spectrum is analytically elusive for general $\mathbf{v}\in\mathbb{S}^{2}$, closed-form expressions are still available in the following cases.
Figure 6: a.) Normalized manipulability and dexterity of MeSCRs with two embedded magnets as functions of normalized magnet position $L_{k_{0}}/L\in[0,1]$: (left) aligned magnetic moments, (right) opposing magnetic moments. b.) Distal end trajectories under various magnet placements $L_{k_{0}}/L$ and moments. Color gradient encodes normalized local manipulability.
Figure 7: a.) Optimization results under $2^{2}$ magnetic orientation combinations. The background color maps represent normalized performance values obtained via exhaustive search. Black lines indicate trajectories of the gradient-based optimization. b.) Optimization results across $2^{3}$ magnetic orientations, comparing the gradient-based method with a heuristic method. Bar plots show the configuration at optimized manipulability with convergence curves overlaid.
III-B1 Analytical Example
Consider a MeSCR with two axially embedded magnets actuated by a spatially uniform field $\mathbf{b}_{u}\in\Pi$ with $\mathbf{v}^{\top}\bar{\mathbf{t}}=0$, where the global manipulability is to be maximized [23]. The two magnets have identical dipole strength. Let one magnet be fixed at the distal end ($L_{k_{1}}=L$), while the position of the other magnet $L_{k_{0}}$ serves as the design variable, subject to the boundary constraint $0<L_{k_{0}}<L$. From Corollary 1, the joint variable simplifies to $\boldsymbol{\theta}_{i}=\theta_{i}\mathbf{n}$. In this configuration, the MeSCR becomes kinematically equivalent to a planar $N$-joint revolute open-chain mechanism. Let $\theta_{i}^{j}\triangleq\sum_{l=i}^{j}\theta_{l}$ denote the cumulative joint angle. In the regime of weak magnetic actuation, the global manipulability index has the linearized expression on $\theta_{k_{0}}^{k_{1}}$: $z\approx L_{k_{0}}(L-L_{k_{0}})\left|\theta_{k_{0}}^{k_{1}}\right|$, where the analytical form is derived from the planar 2-joint revolute open-chain mechanism. Similarly, applying the approximate $\mathbf{G}_{m}$ from (23), the immersion Jacobian also admits the linearized form:
$\mathcal{J}_{\boldsymbol{\theta}}\approx$
$\begin{cases}\!M^{3}\!\sqrt{\Lambda_{k_{0}}^{m}(\Lambda_{k_{1}}^{m}-\Lambda_{k_{0}}^{m})(3\Lambda_{k_{0}}^{m}+\Lambda_{k_{1}}^{m})}\left|\theta_{k_{0}}^{k_{1}}\!\right|,\text{if }\bar{\mathbf{m}}_{k_{0}}\!=\!\bar{\mathbf{m}}_{k_{1}},\\ \!M^{3}\!\sqrt{\Lambda_{k_{0}}^{m}}(\Lambda_{k_{1}}^{m}-\Lambda_{k_{0}}^{m})\left|\theta_{k_{0}}^{k_{1}}\!\right|,\text{if }\bar{\mathbf{m}}_{k_{0}}=-\bar{\mathbf{m}}_{k_{1}},\end{cases}$
The cumulative rotation angle can also be approximated using (17) as $\left|\theta_{k_{0}}^{k_{1}}\right|\approx\frac{L-L_{k_{0}}}{EI}\left|\bar{\mathbf{m}}_{k}^{\top}[\mathbf{b}_{u}]_{\times}\mathbf{n}\right|$. Substituting this into the above expressions for $z$ and $\mathcal{J}_{\boldsymbol{\theta}}$, we can factor the dependence on $L_{k_{0}}$ out of the integral in the objective. The structural optimization problem thus reduces to:
$\max_{L_{k_{0}}}\begin{cases}\!L_{k_{0}}^{2}(L-L_{k_{0}})^{4}\sqrt{4L_{k_{0}}^{2}\!+\!(L-L_{k_{0}})^{2}},\mbox{if}\ \bar{\mathbf{m}}_{k_{0}}\!=\!\bar{\mathbf{m}}_{k_{1}},\\ \!L_{k_{0}}^{2}(L-L_{k_{0}})^{5},\mbox{if}\ \bar{\mathbf{m}}_{k_{0}}\!=\!-\bar{\mathbf{m}}_{k_{1}}.\end{cases}$
Then, the optimal magnet placement is obtained by using the first-order necessary condition as
$L_{k_{0}}^{\ast}\approx\begin{cases}0.368L,\ \text{if }\bar{\mathbf{m}}_{k_{0}}=\bar{\mathbf{m}}_{k_{1}},\\ \frac{2}{7}L,\ \text{if }\bar{\mathbf{m}}_{k_{0}}=-\bar{\mathbf{m}}_{k_{1}}.\end{cases}$
The result is entirely determined by the structural geometry, independent of material or actuation parameters, and thus exhibits scale invariance. Interestingly, when $\bar{\mathbf{m}}_{k_{0}}=\bar{\mathbf{m}}_{k_{1}}$, the ratio of optimal placement $L/L_{k_{0}}^{*}\approx 2.718$ is numerically close to Euler’s number, suggesting a geometric correspondence with the logarithmic spiral structures widely employed in soft continuum design.
We also evaluated the kinematic dexterity of the MeSCR through minimizing the distortion as defined in [24]. It is easy to verify that the distortion density satisfies the assumption of Theorem 5. Thus, we deduce that the optimal configuration of embedded magnets has a boundary distribution. We show numerical simulation results of normalized manipulability and dexterity in Fig. 6a When the magnetic moments are aligned (left graph), the configuration that maximizes manipulability exhibits good agreement with the theoretical result. In contrast, the opposing case (right graph) shows a deviation, which may be attributed to stronger nonlinearity in the index. For both cases, the configurations that minimize distortion consistently place the embedded magnet at the distal end, as expected. Fig. 6b illustrates the reachable distal trajectories under varying magnet placements. Although the optimal placements yield high manipulability along the path, their spatial extent remains limited. This apparent disparity underscores the geometric distinction between local controllability and global reachability, while the latter pertains to the image of the equilibrium map over the actuation submanifold. Therefore, any attempt to maximize overall performance must reconcile the trade-off between local dexterity and global coverage, as the two are not generally aligned.
III-B2 Numerical Examples
For MeSCRs with three or more embedded magnets, deriving a closed-form solution to the optimization problem becomes increasingly intractable. In such cases, we resort to a gradient-based optimization scheme, leveraging the analytical gradient expression provided in (22). Fig. 7a illustrates the optimization trajectories for representative cases with four possible direction combinations. The background colormaps depict the globally normalized performance values obtained via exhaustive enumeration. Across all cases, the gradient-based method consistently converges to configurations that match the global optima, demonstrating its reliability and consistency with full search results.
Moreover, the optimized magnet placements for maximizing dexterity closely follow the geometric insights established in Theorem 5. In all observed cases, the optimizer implicitly minimizes the net torque induced by the embedded magnets. Specifically, when adjacent magnets are aligned, the solution tends to separate them spatially to avoid torque amplification, whereas when they are oppositely oriented, the optimizer brings them closer to enable partial cancellation. These outcomes suggest that the dexterity metric inherently penalizes unbalanced torque distributions that restrict the range of achievable actuation directions.
As shown in Fig. 7b, the heuristic and gradient-based optimization methods yield consistent results both in terms of optimal configuration and convergence behaviors. Combining with the performance maps in Fig. 7a, it demonstrates that manipulability is highly sensitive to the spatial layout and directions of the embedded magnets. In some cases, optimal configurations are observed when magnets are distributed farther apart, highlighting the benefit of spatially diverse actuation sources. In comparison to dexterity, manipulability optimization focuses on maximizing the volume of attainable motion directions. As a result, the optimizer may favor configurations that introduce local torque in exchange for broader kinematic responsiveness.
IV Conclusion and Discussion
This paper presents a unified framework for modeling and optimizing the kinematic performance of multi-magnet embedded soft continuum robots. The proposed EPRB model enables a differentiable representation of equilibrium configurations under magnetic actuation, and supports explicit expressions for both the manipulator Jacobian and Hessian in multi-magnet settings. The analysis of equilibrium well-posedness yields sufficient conditions for solution uniqueness and continuity under bounded actuation, offering a means to ensure continuous configuration tracking in quasi-static operations. The upper bound on controllable degrees of freedom, shown to be twice the number of embedded magnets, offers a simple yet practical rule for relating design parameters to task-space controllability.
A key contribution of this paper lies in the development of a differential geometric optimization framework, where classical kinematic measures are reparametrized over the actuation space via equilibrium immersions. This approach accommodates a wide range of performance criteria, including but not limited to manipulability and dexterity, and admits both analytic and gradient-based solution methods. Although demonstrated in the context of MeSCRs, the formulation applies more broadly to continuum structures actuated by distributed torques, and may offer guidance in the design of other soft or compliant mechanisms where actuation and configuration are implicitly coupled.
Appendix A Proofs of Propositions
A-A Proof of Proposition 1
Noticing that
$\|\mathbf{M}_{i}(\boldsymbol{\theta})\mathbf{b}\|$ $=\left\|\sum_{k\in\mathbb{K},k\geq i}\left(\left[\mathbf{m}_{k}\right]_{\times}^{\top}\mathbf{R}_{0}^{i}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}\right)^{\top}\mathbf{b}(\mathbf{p}_{k})\right\|$
$\leq\left\|\mathbf{R}_{0}^{i}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}\right\|\sum_{k\in\mathbb{K},k\geq i}\left\|\left[\mathbf{m}_{k}\right]_{\times}\mathbf{b}(\mathbf{p}_{k})\right\|$
$=\sum_{k\in\mathbb{K},k\geq i}M_{k}B_{k}\triangleq\mathcal{M}_{i},$
it is convenient to obtain:
$\|\mathbf{M}(\boldsymbol{\theta})\mathbf{b}\|$ $=\sqrt{\sum_{i=0}^{N-1}\|\mathbf{M}_{i}(\boldsymbol{\theta})\mathbf{b}\|^{2}}\leq\sqrt{\sum_{i=0}^{N-1}\mathcal{M}_{i}^{2}}\triangleq\mathcal{M},$
and
$\|\mathbf{M}(\boldsymbol{\theta})\|$ $=\sqrt{\sum_{i=0}^{N-1}\|\mathbf{M}_{i}(\boldsymbol{\theta})\|}\leq\sqrt{N_{m}k_{N_{m}-1}}\max_{k}M_{k}\triangleq\mathcal{M}_{0}.$
A-B Proof of Proposition 2
Given that the Hessian matrix is symmetric, it can be expressed equivalently as
$\frac{\partial^{2}\mathcal{E}_{m}}{\partial\boldsymbol{\theta}\partial\boldsymbol{\theta}^{\top}}=\sum_{k\in\mathbb{K}}\operatorname{sym}\left(\mathbf{D}_{k}^{\top}\mathbf{C}_{k}\mathbf{D}_{k}+\mathbf{Q}^{k}\right)\triangleq\mathbf{S}_{m}.$
Whence,
$\mathbf{D}_{k}=\operatorname{blkdiag}\{\mathbf{R}_{0}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{0}},\dots,\mathbf{R}_{0}^{k}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k}},\underbrace{\mathbf{0}_{3\times 3},\dots,\mathbf{0}_{3\times 3}}_{N-k-1}\},$ (24)
$\mathbf{Q}^{k}=\operatorname{blkdiag}\{\mathbf{Q}_{0},\dots,\mathbf{Q}_{k},\underbrace{\mathbf{0}_{3\times 3},\dots,\mathbf{0}_{3\times 3}}_{N-k-1}\},$
and
$\mathbf{C}_{k}=\operatorname{blkdiag}\{2\mathbf{U}_{k\times k}-\mathbf{I}_{k\times k},\underbrace{0\dots 0}_{N-k-1}\}\otimes\left([\mathbf{b}_{k}]_{\times}^{\top}[\mathbf{m}_{k}]_{\times}\right)$
with $\mathbf{U}_{k\times k}$ denotes the upper triangle matrix with all elements equal to $1$ and $\otimes$ denotes the Kronecker product. Then, using the triangle inequality and sub-multiplicative property, the following inequality holds for the spectrum norm of the Hessian matrix:
$\left\|\mathbf{S}_{m}\right\|$ $=\left\|\sum_{k\in\mathbb{K}}\operatorname{sym}\left\{\mathbf{D}_{k}^{\top}\mathbf{C}_{k}\mathbf{D}_{k}+\mathbf{Q}^{k}\right\}\right\|$ (25)
$\leq\sum_{k\in\mathbb{K}}\left\|\mathbf{D}_{k}^{\top}\right\|\left\|\mathbf{C}_{k}\right\|\left\|\mathbf{D}_{k}\right\|+\left\|\mathbf{Q}^{k}\right\|$
$=\sum_{k\in\mathbb{K}}\left\|\mathbf{C}_{k}\right\|+\left\|\mathbf{Q}^{k}\right\|$
where we used the fact that $\left\|\mathbf{D}_{k}^{\top}\right\|=\left\|\mathbf{D}_{k}\right\|=1$. The spectrum norm of $\mathbf{C}_{k}$ is straightforward to be given by
$\left\|\mathbf{C}_{k}\right\|$ $=\left\|2\mathbf{U}_{k\times k}-\mathbf{I}_{k\times k}\right\|\left\|[\mathbf{b}_{k}]_{\times}^{\top}[\mathbf{m}_{k}]_{\times}\right\|$ (26)
$\leq\frac{4k}{\pi}\|\mathbf{b}_{k}\|\|\bar{\mathbf{m}}_{k}\|=\frac{4k}{\pi}M_{k}B_{k}$
For small rotation angles, we note that
$\lim_{\boldsymbol{\theta}\to\mathbf{0}}\left\|\mathbf{Q}^{k}\right\|$ $=\lim_{\boldsymbol{\theta}\to\mathbf{0}}\max_{i=0,\dots,k}\left\{\left\|\mathbf{Q}_{i}\right\|\right\}=0.$ (27)
Combing (25), (26), and (27), the Jacobian of the vector-valued function $\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$ satisfies:
$\left\|\frac{\partial}{\partial\boldsymbol{\theta}}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}\right\|=\left\|-\mathbf{S}_{m}\right\|\leq\sum_{k\in\mathbb{K}}\frac{4k}{\pi}M_{k}B_{k}\triangleq\mathcal{L}.$
Therefore, the function is Lipschitz continuous on $\mathbb{R}^{3N}$ with the Lipschitz constant $\mathcal{L}$.

### v2

@@
-Unified Modeling and Structural Optimization of Multi-magnet Embedded Soft Continuum Robots for Enhanced Kinematic Performances
+Enhancing Kinematic Performances of Soft Continuum Robots for Magnetic Actuation
 Abstract
-This paper presents a unified modeling and optimization framework to enhance the kinematic performance of multi-magnet embedded soft continuum robots (MeSCRs).
-To this end, we establish a differentiable system formulation based on an extended pseudo-rigid-body model.
-This formulation enables analysis of the equilibrium well-posedness and the geometry of the induced configuration under magnetic actuation.
-In particular, we show that the maximum controllable degrees of freedom of a MeSCR equal twice the number of embedded magnets.
-We subsequently develop a structural optimization framework based on differential geometry that links classical kinematic measures (e.g., manipulability and dexterity) to the configuration of embedded magnets.
-The resulting optimization condition reveals that improving local performance requires structurally modulating the spectrum of the configuration space metric to counteract its distortion.
-Closed-form solutions for optimal magnet configurations are derived under representative conditions, and a gradient-based numerical method is proposed for general design scenarios.
-Simulation studies validate the effectiveness of the proposed framework.
-Index Terms:
-Magnetic soft continuum robots, pseudo-rigid-body modeling, kinematic performances, structure optimization.
+Soft continuum robots achieve complex deformation through elastic equilibrium, making their reachable motions governed jointly by structural design and actuation-induced mechanics.
+This work develops a general formulation that integrates equilibrium computation with kinematic performances by evaluating Riemannian Jacobian spectra on the equilibrium manifold shaped by internal/external loading.
+The resulting framework yields a global performance functional that directly links structural parameters, actuation inputs, and the induced configuration space geometry.
+We apply this general framework to magnetic actuation.
+Analytical characterization is obtained under weak uniform fields, revealing optimal placement and orientation of the embedded magnet with invariant scale properties.
+To address nonlinear deformation and spatially varying fields, a two-level optimization algorithm is developed that alternates between energy based equilibrium search and gradient based structural updates.
+Simulations and physical experiments across uniform field, dipole field, and multi-magnet configurations demonstrate consistent structural tendencies: aligned moments favor distal or mid-distal solutions through constructive torque amplification, whereas opposing moments compress optimal designs toward proximal regions due to intrinsic cancellation zones.
 I Introduction
-Magnetically actuated soft continuum robots have demonstrated significant potential in medical procedures such as vascular intervention and gastrointestinal examination[1].
-Early designs typically embedded a single magnet at the distal end to enable basic magnetic manipulation[2, 3, 4].
-However, the limited torque generated by such configurations restricts the controllable degrees of freedom (DoF) less than two[5], significantly constraining spatial maneuverability.
-Recent advances in magnetic materials and microfabrication have enabled the integration of multiple micro-magnets along the continuum body[6, 7], giving rise to multi-magnet embedded soft continuum robots (MeSCRs).
-These systems provide multiple distributed sources of actuation torque, substantially enhancing kinematic performance and expanding the design space.
-Various strategies have been proposed to exploit these capabilities, such as using movable magnets to enlarge reachable workspace[8], designing magnet layouts for multimodal deformation[9, 10], or inducing axial twisting via circumferential arrangements[11].
-These efforts can be seen as structural design choices, analogous to tuning joint orientations or link lengths in rigid manipulators to improve kinematic capabilities.
-Despite this progress, the absence of a unified modeling framework for MeSCRs remains a major barrier.
-Unlike rigid-link manipulators, which benefit from well-established analytic tools for performance analysis and optimization, soft continuum robots lack a theoretical foundation that systematically links structure to kinematic function[12].
-As a result, many mature strategies developed in rigid manipulators cannot be directly applied to MeSCRs, motivating the need for a new modeling paradigm that bridges this gap.
-Current MeSCR modeling approaches typically fall into two categories: Cosserat rod models and pseudo-rigid-body (PRB) models.
-The Cosserat formulation provides high-fidelity representations and has been integrated with magnetic dipole actuation models for motion control[13, 14].
-However, it requires solving boundary value problems derived from variational principles, making it analytically intractable and computationally intensive for iterative design.
-By contrast, the PRB model discretizes flexible structures into rigid links with elastic joints[15, 16], yielding a Lagrangian formulation analogous to rigid manipulators.
-This enables unified expressions for forward kinematics and simplifies equilibrium computation via unconstrained finite-dimensional optimization[17].
-Early applications in single-magnet MeSCRs have demonstrated effective Jacobian-based control strategies[18, 19], and rank-based analysis has shown that such systems can control at most two DoF[5].
-Recent work has extended the PRB formulation to multi-magnet settings[20, 10], preliminarily validating its capacity to simulate complex deformations.
-Nevertheless, the PRB model also suffers from several inherent limitations.
-The choice of joint discretization entails a trade-off between modeling fidelity and computational cost.
-In particular, due to the lack of theoretical analysis on well-posedness, existing methods typically use numerical methods to analyze the multiplicity and jump phenomena of solutions[21].
-Moreover, to our best knowledge, complete closed-form expressions for Jacobian (simplified expression can be found in [20, 19]) and Hessian matrices under multi-magnet configurations have not been established.
-To address the limitations outlined above, the first part of this paper investigates an extended-pseudo-rigid-body (EPRB) model[17, 19], which introduces serial spherical joints to enable unified modeling of spatial bending and torsional responses.
-The EPRB model possesses a generalized expressive capacity: when the elastic modulus approaches zero, it naturally degenerates into a classical rigid-link manipulator model.
-Conversely, as the number of embedded magnets tends to infinity, it approximates the behavior of a magnetic continuum[22].
-Based on this formulation, we conduct a well-posedness analysis of the static equilibrium equations, encompassing uniqueness, boundedness, and convergence properties.
-Furthermore, we derive complete analytical expressions for the Jacobian and Hessian matrices under multi-magnet configurations, thereby providing theoretical support for both kinematic analysis and controller design.
-Building upon the proposed EPRB model, the second part of this paper further demonstrates its ability to facilitate geometric analysis and structural optimization of soft continuum robots.
-We begin by rigorously proving a key structural result: the maximum controllable degrees of freedom in a magnetically actuated soft continuum robot is exactly twice the number of embedded magnets.
-This finding corroborates prior empirical observations reported in many existing literature[5, 20].
-Additionally, we show that when magnets are arranged axially, the EPRB model naturally reduces to the classical PRB formulation, and a closed-form solution to the equilibrium configuration can be obtained under small-angle premises.
-As structural optimization often aims to enhance a robot’s kinematic performance, we draw inspiration from rigid manipulator design, where Jacobian-based kinematic measures (e.g., manipulability and dexterity) are widely employed to guide structural synthesis and task planning[23, 24].
-We develop a differential geometry-based optimization framework around kinematic measures, which connects kinematic performances to the configuration of embedded magnets in MeSCR.
-Unlike rigid-link manipulators, whose structural and performance optimization can be directly formulated in joint space, soft continuum robots exhibit implicitly coupled configurations governed by elastomagnetic equilibrium.
-As such, optimization must be performed over the actuation or design space, with performance indices reparametrized through the configuration manifold.
-To resolve this, we reinterpret the equilibrium set as a smooth submanifold and apply tools from geometric measure theory to reparametrize global performance objectives over the actuation space.
-We derive closed-form optimal magnet configuration in representative cases and develop an efficient gradient-based numerical method for the general scenario.
+Soft continuum robots have gained growing attention for tasks involving compliant interaction, dexterous access, and safe manipulation in complex or confined environments.
+Their ability to realize smooth, multi-segment deformation without rigid joints supports applications in minimally invasive navigation, inspection, and human-centered tasks.
+Recent advancements have significantly improved perception, control, and structural adaptability in soft continuum systems.
+Examples include proprioceptive designs capable of simultaneous shape estimation and contact detection [1], learning based deformation estimation under multi-point interaction [2], optimal control strategies for rotation-coupled bending [3], and variable curvature mechanisms that enhance configurability [4].
+These developments collectively broaden the functional space of soft continuum manipulation and motivate new models that can capture their nonlinear, compliant behavior.
+Magnetic actuation has emerged as a particularly promising approach for soft continuum robots.
+Compared with tendon-driven or pneumatic mechanisms, magnetic actuation allows wireless control, allows miniaturization to submillimeter scales, and reduces onboard complexity.
+Recent studies demonstrate substantial gains in workspace, dexterity, and navigation capability across diverse magnetic architectures.
+Robots with distributed magnetization patterns have shown enhanced deformation diversity [5], movable opposite magnets have enabled workspace expansion [6], and rotating field-assisted designs have achieved buckling free insertion in tortuous lumens [7].
+High dexterity helical magnetic robots further extend steerability in endovascular navigation [8].
+In parallel, modeling frameworks for magnetic and elastic coupling have evolved rapidly.
+Analytical and semi-analytical models have been introduced for large deformation magnetoelastic behavior, including kinetostatic formulations capable of capturing multistage bending under complex dipole fields [9], modular magnetization profile evaluation [10], and optimized structural patterns for enhanced deformability.
+Additional works explore interactions between magneto elasticity, external loading, and nonlinear geometries, providing foundations for real-time control and structural tuning.
+Together, these contributions highlight the importance of accurate equilibrium modeling and structural design for improving kinematic performance.
+However, several core challenges remain unresolved.
+Magnetic soft continuum robots do not operate on a fixed kinematic map.
+Their attainable configurations are determined by the equilibrium manifold produced by the interaction between elastic restoring forces and magnetic loading.
+This manifold is highly nonlinear and sensitive to magnet placement, field variation, torque cancellation, and bifurcation events.
+Existing design methods often rely on task-specific heuristics [6], empirical tuning of magnet distributions [5], data-driven control [11, 12], or simplified kinematic approximations that overlook the intrinsic coupling between mechanics and actuation.
+Learning-based or adaptive controllers [13, 1] enhance robustness, yet they offer limited insight into the structural sources underlying the equilibrium geometry.
+Moreover, commonly used performance indices such as manipulability and dexterity, being defined in Euclidean coordinates, may only partially capture the curvature of the equilibrium manifold under magnetic loading.
+As a result, structural optimization and performance evaluation are often decoupled, limiting the ability to systematically ensure actuation feasibility, stability, and kinematic quality.
+These gaps motivate the need for a unified formulation that couples equilibrium with kinematic performance.
+Since the robot’s motion is realized only through equilibrium configurations, performance should be evaluated on the manifold shaped by internal and external loading.
+It further requires a structural optimization scheme capable of handling nonlinear magnetoelastic interactions, diverse field geometries, and multi-magnet architectures recently explored in continuum and catheter-like robots [14, 6, 7].
+In this study, we develop a unified framework and demonstrate its theoretical and practical utility across homogeneous field and dipole field actuation.
+We leverage Riemannian Jacobian spectra to characterize manipulability and dexterity on the equilibrium manifold and integrate these indices with an energy-based equilibrium computation.
+This formulation enables physically consistent performance evaluation and reveals structural trends that persist across actuation regimes, including robots with modular magnetization segments [10], multi-magnet bodies [5], and clinically oriented continuum tools [15, 16].
+The main contributions of this work are summarized as follows:
+  1. 1.
+    We establish a unified structural optimization framework that couples magnetic energy-based equilibrium analysis with Riemannian Jacobian spectral evaluation on the equilibrium manifold.
+By formulating the equilibrium set as an actuation-dependent submanifold and computing kinematic performance on its pullback metric.
+  2. 2.
+    We derive analytical results for soft continuum robots actuated by weak uniform magnetic fields, including closed-form optimality conditions for magnet placement and orientation.
+These theoretical characterizations reveal scale-invariant structural principles that are further validated through numerical optimization.
+  3. 3.
+    We extend the proposed framework to general dipole-generated magnetic fields and apply the full two-level optimization scheme to multi-magnet configurations.
+Comprehensive numerical studies and corresponding physical experiments demonstrate consistent trends in equilibrium behavior and kinematic performance across diverse actuation paths.
+The remainder of this paper is organized as follows.
+Sec.
+II presents the modeling framework, including the pseudo-rigid-body representation, magnetic torque formulation, and energy-based equilibrium analysis.
+Sec.
+III defines the kinematic performance metrics and formulates the structural optimization problem.
+Sec.
+IV introduces the analytical and numerical optimization methods.
+Sec.
+V provides case studies and numerical validations under uniform and general magnetic fields.
+Finally, Sec.
+VI discusses the implications of the proposed framework and concludes the paper.
 Throughout the technical description, vectors and matrices are defined using lowercase and uppercase bold symbols, respectively, whereas scalars are not bolded.
-II Modeling
-This section introduces the extended pseudo-rigid-body model to describe the MeSCR.
-Complete closed-form expressions for the manipulator Jacobian and Hessian matrices are derived.
-We further investigate the well-posedness of the equilibrium solution.
-These results provide the necessary mathematical foundation for subsequent kinematic analysis and performance optimization.
-II-A Forward Kinematics and Manipulator Jacobian
-Figure 1: Illustration of the basic concept of the EPRB model on describing the MeSCR.
-The continuous body is discretized by a series of rotational rigid rods with joints and frames attached.
-The bending and twisting deformations of the MeSCR are described by three joint variables $\theta_{\mathbf{t}}^{i}$, $\theta_{\mathbf{u}}^{i}$, and $\theta_{\mathbf{v}}^{i}$.
-To meet the demands of entering narrow spaces and achieving complex deformations, MeSCRs are typically designed with slender structures.
+II Kinematics of Soft Continuum Robots
+II-A Kinematic Mapping
+Figure 1: Illustration of the basic concept of the PRB model on describing the SCR.
+The soft body is discretized into rigid rods connected by joints with local frames.
+The configuration is represented by the joint variables $\theta_{\mathbf{t}}^{i}$, $\theta_{\mathbf{u}}^{i}$, and $\theta_{\mathbf{v}}^{i}$.
+The system reaches an equilibrium state (②) when the torques of the actuators (①) are balanced by the elastic restoring moments, and the joint deformation is stored as elastic strain.
+We consider SCRs with slender structures, whose length is significantly greater than their radius.
 Certain well-founded assumptions contribute to simplifying the model [17], such as the rod remaining free of shear deformation and its cross-section always being orthogonal to the centerline.
@@
 In contrast, we assume only that the continuum is inextensible, with twisting characterized by extended variables [19].
-Within the EPRB model framework, a continuum robot of length $L$ is described by a series of rigid rods.
+Within the PRB model framework, the SCR with length $L$ is described by a series of rigid rods.
 As shown in Fig.
-1, $N$ spherical joints are distributed along the centerline of the MeSCR at points of interest, connected by $N$ rigid rods with lengths $l_{i},i=0,\dots,N-1$.
+1, $N$ spherical joints are distributed along the centerline of the SCR at points of interest, connected by $N$ rigid rods with lengths $l_{i},i=0,\dots,N-1$.
 We use the Voronoi region of the joint to scale the potential energies, which is defined as $\ell_{i}=(l_{i-1}+l_{i})/2,i=0,\dots,N-1$.
-An orthogonal local frame $\{\bar{\mathbf{t}}_{i},\bar{\mathbf{u}}_{i},\bar{\mathbf{v}}_{i}=\bar{\mathbf{t}}_{i}\times\bar{\mathbf{u}}_{i}\}$ is attached to each joint, with overbars indicating that the MeSCR is in a straightened reference state, and the tangent vector $\bar{\mathbf{t}}_{i}$ remains aligned with the direction of the successive rod.
+An orthogonal local frame $\{\bar{\mathbf{t}}_{i},\bar{\mathbf{u}}_{i},\bar{\mathbf{v}}_{i}=\bar{\mathbf{t}}_{i}\times\bar{\mathbf{u}}_{i}\}$ is attached to each joint, with overbars indicating that the SCR is in a straightenedl reference state, and the tangent vector $\bar{\mathbf{t}}_{i}$ remains aligned with the direction of the successive rod.
 The local frame is used to determine the rotation axis for each joint, expressed as $\boldsymbol{\theta}_{i}=\theta_{\mathbf{t}}^{i}\bar{\mathbf{t}}+\theta_{\mathbf{u}}^{i}\bar{\mathbf{u}}+\theta_{\mathbf{v}}^{i}\bar{\mathbf{v}}\in\mathbb{R}^{3}$, where $\theta_{\mathbf{t}}^{i},\theta_{\mathbf{v}}^{i},\theta_{\mathbf{u}}^{i}\in\mathbb{R}$ are three independent joint variables.
 The associated rotation matrix is defined as $\mathbf{R}_{i}(\boldsymbol{\theta}_{i})=\mathbf{e}^{[\boldsymbol{\theta}_{i}]_{\times}}\in\mathrm{SO}(3)$, and we define a useful shorthand notation $(\mathbf{R}_{i}^{j})$ to represent the product of rotation matrices that $\mathbf{R}_{i}^{j}=\mathbf{R}_{i}\mathbf{R}_{i+1}\cdots\mathbf{R}_{j}$.
-With the notation, local frame axes in a deformed state (②), such as the tangent vector, can be easily computed by $\bar{\mathbf{t}}=\mathbf{R}_{0}^{i}\bar{\mathbf{t}}_{i}$ from the reference state (①).
+With the notation, local frame axes in a deformed state (②), such as the tangent vector, can be easily computed by $\mathbf{t}_{i}=\mathbf{R}_{0}^{i}\bar{\mathbf{t}}_{i}$ from the reference state (①).
 The resulting elastic strain on the $i$th joint will be used to compute the elastic potential energy later.
-The configuration space $(\boldsymbol{\theta})$ of the MeSCR can be written as
-$\boldsymbol{\theta}=\left[\boldsymbol{\theta}_{0},\dots,\boldsymbol{\theta}_{N-1}\right]^{\top}\in\mathbb{T}\subset\mathbb{R}^{3N_{m}}.$ (1)
-Let a fixed base frame $\{s\}$ attach to joint 00, frame $\{b\}$ attach to the distal end, and ${}_{s}^{b}\bar{\mathbf{H}}\in\mathrm{SE}(3)$ be the associated distal pose when the MeSCR is in its zero position.
-The screw axes in the fixed base frame are of the form:
-$\boldsymbol{\xi}_{i}=\begin{bmatrix}\boldsymbol{\theta}_{i}\\ -\boldsymbol{\theta}_{i}\times\bar{\mathbf{p}}_{i}\end{bmatrix}$
-where $\bar{\mathbf{p}}_{i}$ is the reference position of the $i$th joint associated with $l_{i}$.
-Then, the forward kinematics of the MeSCR can be represented in the following product of exponentials:
-${}_{s}^{b}\mathbf{H}(\boldsymbol{\theta})=\mathbf{e}^{\boldsymbol{\xi}_{0}^{\wedge}}\mathbf{e}^{\boldsymbol{\xi}_{1}^{\wedge}}\cdots\mathbf{e}^{\boldsymbol{\xi}_{N-1}^{\wedge}}{}_{s}^{b}\bar{\mathbf{H}}(\mathbf{0})\in\mathrm{SE}(3).$
+The configuration space $(\boldsymbol{\theta})$ of the SCR can be written as
+$\boldsymbol{\theta}=\left[\boldsymbol{\theta}_{0},\dots,\boldsymbol{\theta}_{N-1}\right]^{\top}\in\mathbb{T}\subset\mathbb{R}^{3N}.$ (1)
+Let a fixed base frame $\{s\}$ attach to joint 0, frame $\{b\}$ attach to the distal end, and $_{b}^{s}\bar{\mathbf{H}}\in\mathrm{SE}(3)$ be the associated distal pose when the SCR is in its zero position.
+The screw axes in the fixed base frame are of the form
+$\boldsymbol{\xi}_{i}=\begin{bmatrix}\boldsymbol{\theta}_{i}\\ -\boldsymbol{\theta}_{i}\times\bar{\mathbf{p}}_{i}\end{bmatrix}\in\mathbb{R}^{6}$
+where $\bar{\mathbf{p}}_{i}$ is the reference position of the $i$th joint associated with $l_{i}$. Then, the forward kinematics of the SCR can be represented in the following product of exponentials:
+$_{b}^{s}\mathbf{H}(\boldsymbol{\theta})=\mathbf{e}^{\boldsymbol{\xi}_{0}^{\wedge}}\mathbf{e}^{\boldsymbol{\xi}_{1}^{\wedge}}\cdots\mathbf{e}^{\boldsymbol{\xi}_{N-1}^{\wedge}}{}_{b}^{s}\bar{\mathbf{H}}(\mathbf{0})\in\mathrm{SE}(3).$ (2)
 We next present the analytical expression of the space Jacobian $\mathbf{J}_{\boldsymbol{\theta}}(\boldsymbol{\theta})\in\mathbb{R}^{6\times 3N}$, defined in fixed frame coordinates.
-The space Jacobian relates small changes in joint space to distal pose as $\delta\mathbf{h}=\mathbf{J}_{\boldsymbol{\theta}}\delta\boldsymbol{\theta}$, where $\mathbf{h}=\mathrm{Log}\left({}_{s}^{b}\mathbf{H}\right)\in\mathbb{R}^{6}$ represents the distal pose in Cartesian coordinates.
-To facilitate calculation, the Jacobian matrix is partitioned into blocks, with each block corresponding to a joint space group as
+The space Jacobian relates small changes in joint space to distal pose as $\delta\mathbf{h}=\mathbf{J}_{\boldsymbol{\theta}}\delta\boldsymbol{\theta}$, where $\mathbf{h}=\mathrm{Log}\left({}_{b}^{s}\mathbf{H}\right)\in\mathbb{R}^{6}$ represents the distal pose in Cartesian coordinates.
+The Jacobian matrix is partitioned into blocks, with each block corresponding to a joint space group:
 $\mathbf{J}_{\boldsymbol{\theta}}^{s}=\left[\frac{\partial\mathbf{h}}{\partial\boldsymbol{\theta}_{0}},\ \frac{\partial\mathbf{h}}{\partial\boldsymbol{\theta}_{1}},\ \cdots\ ,\frac{\partial\mathbf{h}}{\partial\boldsymbol{\theta}_{N-1}}\right]$ (3)
@@
 $\frac{\partial\mathbf{h}}{\partial\boldsymbol{\theta}_{i}}=\mathbf{Ad}_{\mathbf{e}^{\boldsymbol{\xi}_{0}^{\wedge}}\cdots\mathbf{e}^{\boldsymbol{\xi}_{i-1}^{\wedge}}}\mathbf{J}_{\mathbf{H}_{l}}^{\boldsymbol{\xi}_{i}}\frac{\partial\boldsymbol{\xi}_{i}}{\partial\boldsymbol{\theta}_{i}}\in\mathbb{R}^{6\times 3}$
-with $\frac{\partial\boldsymbol{\xi}_{i}}{\partial\boldsymbol{\theta}_{i}}=\begin{bmatrix}\mathbf{I}_{3},\ -[\bar{\mathbf{p}}_{i}]_{\times}\end{bmatrix}^{\top}$, and $\mathbf{J}_{\mathbf{H}_{l}}^{\boldsymbol{\xi}_{i}}$ is the left Jacobian of the $\mathrm{SE}(3)$ manifold, as discussed in [27].
+with $\frac{\partial\boldsymbol{\xi}_{i}}{\partial\boldsymbol{\theta}_{i}}=\begin{bmatrix}\mathbf{I}_{3},\ -[\bar{\mathbf{p}}_{i}]_{\times}\end{bmatrix}^{\top}$, and $\mathbf{J}_{\mathbf{H}_{l}}^{\boldsymbol{\xi}_{i}}$ is the left Jacobian of the $\mathrm{SE}(3)$ manifold as discussed in [20].
 To conclude, we note the following relationship between the space Jacobian and the body Jacobian $(\mathbf{J}_{\boldsymbol{\theta}}^{b}(\boldsymbol{\theta}))$ in case of need:
 $\mathbf{J}_{\boldsymbol{\theta}}^{b}=\mathbf{Ad}_{\mathbf{e}^{\boldsymbol{\xi}_{0}^{\wedge}}\cdots\mathbf{e}^{\boldsymbol{\xi}_{N-1}^{\wedge}}}^{-1}\mathbf{J}_{\boldsymbol{\theta}}^{s}.$ (4)
-Also for brevity, $\mathbf{J}_{\boldsymbol{\theta}}$ will be referred to $\mathbf{J}_{\boldsymbol{\theta}}^{s}$ unless specified.
+Also for brevity, $\mathbf{J}_{\boldsymbol{\theta}}$ will be referred to $\mathbf{J}_{\boldsymbol{\theta}}^{b}$ unless specified.
 II-B Equations of Motion
-Figure 2: Illustration of a quasi-static state of the MeSCR under an external magnetic field $\mathbf{b}(\mathbf{p}_{k})$.
-Multiple embedded micromagnets at corresponding joints generate magnetic torques $\mathbf{m}_{k}\times\mathbf{b}_{k}$ that induce deformation in the MeSCR.
-When the magnetic torque and elastic moments reach the equilibrium, the MeSCR attains the quasi-static state.
-The MeSCR’s equations of motion are derived from the Lagrangian mechanics, which are formulated based on the MeSCR’s potential energy.
-As shown in Fig.
-2, the MeSCR is allowed to bend, twist, and magnetic actuatable.
-Thus, the total potential energy features elastic strains and magnetic moments, which are decomposed into the elastic $(\mathcal{E}_{e})$ and magnetic $(\mathcal{E}_{m})$ potential energies as
-$\mathcal{E}=\mathcal{E}_{e}+\mathcal{E}_{m}.$ (4)
-The respective energies are formulated as follows.
-II-B1 Elastic Potential Energy
-In general deformation conditions, the elastic strain can be approximated as a linear function of the joint angle $\boldsymbol{\theta}_{i}$, [15, 28, 26].
-The quadratic form of the elastic strain formulates the elastic potential energy as
-$\mathcal{E}_{e}$ $=\frac{1}{2}\sum_{i=0}^{N-1}(\boldsymbol{\theta}_{i}-\bar{\boldsymbol{\theta}}_{i})^{\top}\boldsymbol{\Lambda}_{i}\boldsymbol{(}\boldsymbol{\theta}_{i}-\bar{\boldsymbol{\theta}}_{i})$
-$=\frac{1}{2}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})^{\top}\boldsymbol{\Lambda}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})$
-where $\boldsymbol{\Lambda}_{i}=\operatorname{diag}(\frac{2G_{i}I_{i}}{\ell_{i}},\frac{E_{i}I_{i}}{\ell_{i}},\frac{E_{i}I_{i}}{\ell_{i}})$, $G_{i}$, $E_{i}$, $I_{i}$ are the shear modulus, Young’s modulus, and cross-sectional inertia of the $i$th joint, respectively.
-The block-diagonal matrix $\boldsymbol{\Lambda}$ is composed of $\boldsymbol{\Lambda}_{i}$, referred to as the joint stiffness matrix.
-We also note $G_{i}=\frac{E_{i}}{2(1+\rho_{i})}$ for the isotropic material used in the MeSCR, where $\rho_{i}$ represents the Poisson’s ratio.
-The material properties of each joint are determined by whether it comprises a hard-magnetic material or a soft matrix, both of which exhibit a positive Poisson’s ratio.
-II-B2 Magnetic Potential Energy
-As shown in Fig.
-2, each embedded micromagnet is treated as an independent joint, with additional joints added on either side to assist the description.
+The equations of motion are derived within the framework of Lagrangian mechanics, formulated in terms of the potential energies stored in the SCR.
+The SCR is assumed to undergo bending and twisting deformations and to be driven by generalized actuation sources, which can represent either distributed or localized internal/external torques.
+Accordingly, the total potential energy is composed of elastic strains and actuation-induced moments, decomposed into the elastic $(\mathcal{E}_{e})$ and actuation $(\mathcal{E}_{u})$ energies:
+$\mathcal{E}(\boldsymbol{\theta},\boldsymbol{u})=\mathcal{E}_{e}(\boldsymbol{\theta})+\mathcal{E}_{u}(\boldsymbol{\theta},\boldsymbol{u}),$ (5)
+where $\boldsymbol{u}$ denotes the actuation input.
+For general deformations, the elastic strain at each joint can be approximated as a linear function of the joint angle $\boldsymbol{\theta}_{i}$ [19, 18].
+The quadratic form of the elastic strain yields the elastic potential energy
+$\mathcal{E}_{e}=\frac{1}{2}\sum_{i=0}^{N-1}(\boldsymbol{\theta}_{i}-\bar{\boldsymbol{\theta}}_{i})^{\top}\boldsymbol{\Lambda}_{i}\boldsymbol{(}\boldsymbol{\theta}_{i}-\bar{\boldsymbol{\theta}}_{i})=\frac{1}{2}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})^{\top}\boldsymbol{\Lambda}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})$
+where $\boldsymbol{\Lambda}_{i}=\operatorname{diag}\left(\tfrac{2G_{i}I_{i}}{\ell_{i}},\tfrac{E_{i}I_{i}}{\ell_{i}},\tfrac{E_{i}I_{i}}{\ell_{i}}\right)$, with $G_{i}$, $E_{i}$, and $I_{i}$ denoting the shear modulus, Young’s modulus, and cross-sectional inertia of the $i$-th joint, respectively.
+The block-diagonal matrix $\boldsymbol{\Lambda}=\operatorname{blkdiag}(\boldsymbol{\Lambda}_{0},\cdots,\boldsymbol{\Lambda}_{N-1})$ is referred to as the stiffness matrix.
+For isotropic material, $G_{i}=\tfrac{E_{i}}{2(1+\rho_{i})}$, with $\rho_{i}$ being the Poisson’s ratio.
+The actuation potential energy $\mathcal{E}_{u}(\boldsymbol{\theta},\boldsymbol{u})$ is left in a general form, representing the work done by internal or external actuators.
+Its gradient with respect to $\boldsymbol{\theta}$ produces the generalized actuation torques $\nabla_{\boldsymbol{\theta}}\mathcal{E}_{u}(\boldsymbol{\theta},\boldsymbol{u})$.
+By Lagrangian formulation, neglecting inertial terms in the quasi-static regime, the equilibrium configuration is determined by the equilibrium condition:
+$\nabla_{\boldsymbol{\theta}}\mathcal{E}_{e}(\boldsymbol{\theta})+\nabla_{\boldsymbol{\theta}}\mathcal{E}_{u}(\boldsymbol{\theta},\boldsymbol{u})=\boldsymbol{\tau}(\boldsymbol{\theta}),$ (6)
+where $\boldsymbol{\tau}(\boldsymbol{\theta})$ denotes additional generalized torques such as gravity or contact forces.
+II-C Kinematic Performances
+We next introduce the definition of kinematic performance of SCRs with coordinate-invariance.
+We begin by recalling that a Riemannian manifold is a smooth manifold equipped with an inner product on each tangent space that varies smoothly with the base point.
+The associated volume form provides a canonical measure for integration.
+Let $\mathbb{T}\subset\mathbb{R}^{n}$ be the manifold and use the Euclidean metric, then the induced volume form reduces to the standard Lebesgue measure $\Omega_{\mathbb{T}}=\mathrm{d}\boldsymbol{\theta}$.
+Let $^{s}_{b}\mathbf{H}:\mathbb{T}\to SE(3)$ be the forward kinematics of the SCR and $\mathbf{J}_{\boldsymbol{\theta}}$ its geometric Jacobian.
+Equipping $SE(3)$ with the Euclidean metric induces the pullback metric $(\mathbf{J}_{\boldsymbol{\theta}}^{\top}\mathbf{J}_{\boldsymbol{\theta}})$, whose nonzero eigenvalues $\{\lambda_{i}(\mathbf{J}_{\boldsymbol{\theta}}^{\top}\mathbf{J}_{\boldsymbol{\theta}})\}$ characterize the local distortion of $^{s}_{b}\mathbf{H}$.
+Any symmetric function of these eigenvalues defines a spectral functional $z(\lambda_{1},\ldots,\lambda_{m})$, which serves as a general mathematical form of a kinematic performance measure.
+A local performance density can then be defined in general as any symmetric spectral functional $z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta}))\triangleq z(\lambda_{1}(\boldsymbol{\theta}),\ldots,\lambda_{m}(\boldsymbol{\theta})),$ with typical choices including
+$z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta}))=\left\{\begin{aligned} &z_{\mathrm{vol}}(\boldsymbol{\theta})=\sqrt{\det(\mathbf{J}_{\boldsymbol{\theta}}\mathbf{J}_{\boldsymbol{\theta}}^{\top})},\\ &z_{\mathrm{dis}}(\boldsymbol{\theta})=\tfrac{1}{2}\operatorname{tr}(\mathbf{J}_{\boldsymbol{\theta}}^{\top}\mathbf{J}_{\boldsymbol{\theta}}),\\ &z_{\kappa}(\boldsymbol{\theta})=\tfrac{\sigma_{\max}(\mathbf{J}_{\boldsymbol{\theta}})}{\sigma_{\min}(\mathbf{J}_{\boldsymbol{\theta}})},\end{aligned}\right$.
+(7)
+corresponding respectively to volume gain, distortion energy, and condition-number-based dexterity [21, 22, 23].
+These local measures are extended to global indices by integration with respect to $\Omega_{\mathbb{T}}$ as $\mathcal{Z}\triangleq\int_{\mathbb{T}}z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta}))\ \Omega_{\mathbb{T}}.$ However, since feasible configurations of the SCR are constrained by the equilibrium condition (6), the effective integration domain is the solution set $\mathbb{T}^{*}=\{\boldsymbol{\theta}\in\mathbb{T}\mid\nabla_{\boldsymbol{\theta}}\mathcal{E}(\boldsymbol{\theta},\boldsymbol{u})=\boldsymbol{\tau}(\boldsymbol{\theta})\}$.
+If the constraint has constant full rank with respect to $\boldsymbol{\theta}$, $\mathbb{T}^{*}$ is a smooth submanifold of $\mathbb{T}$, and the induced Riemannian volume form can be used.
+More generally, including neighborhoods of singularities or branch points, the equilibrium set $\mathbb{T}^{*}(\boldsymbol{u})$ is countably rectifiable with intrinsic dimension $\dim(\mathbb{T}^{*}(\boldsymbol{u}))$, hence it carries a well-defined Hausdorff measure $\mathcal{H}^{\dim(\mathbb{T}^{*})}$ and supports geometric integration.
+Consequently, the constrained global performance can be written as an iterated integral over the actuation parameter space,
+$\mathcal{Z}=\int_{\mathbb{U}}\bigg(\int_{\mathbb{T}^{*}}z\!\left({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta})\right)\,\mathrm{d}\mathcal{H}^{\dim(\mathbb{T}^{*})}(\boldsymbol{\theta})\bigg)\,\Omega_{\mathbb{U}},$
+where $\Omega_{\mathbb{U}}$ denotes the volume measure on the admissible actuation domain $\mathbb{U}$.
+III Structural Optimization Methodology
+III-A Problem Formulation
+Let $\mathbf{a}\in\mathbb{A}$ denote the vector of structural design parameters, such as segment lengths, joint stiffness blocks, or orientations of embedded units.
+The admissible set $\mathbb{A}$ is specified by manufacturing limits and safety considerations.
+For each candidate design $\mathbf{a}$, the forward kinematics $^{s}_{b}\mathbf{H}(\boldsymbol{\theta};\mathbf{a})$ and the corresponding Jacobian $\mathbf{J}_{\boldsymbol{\theta}}(\boldsymbol{\theta};\mathbf{a})$ are uniquely determined, and hence the local performance density $z$ and its global counterpart $\mathcal{Z}(\mathbf{a})$ can be evaluated using (7).
+The structural optimization problem is then formulated as
+$\min_{\mathbf{a}}$ $\mathcal{Z}(\mathbf{a})\;\triangleq\;\int_{\mathbb{T}}z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta};\mathbf{a}))\,\Omega_{\mathbb{T}}$ (8)
+[…]
+where $\mathcal{A}(\mathbf{a})\leq 0$ collects algebraic feasibility constraints.
+Consequently, the global index associated with the structural parameter $(\mathbf{a})$ can be equivalently written as an iterated integral over the actuation domain and the corresponding solution manifold:
+$\min_{\mathbf{a}}\mathcal{Z}(\mathbf{a})$ $=\int_{\mathbb{U}}\Bigg(\int_{\mathbb{T}^{*}}z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta};\mathbf{a}))\,\mathrm{d}\mathcal{H}^{\dim(\mathbb{T}^{*})}(\boldsymbol{\theta})\Bigg)\,\Omega_{\mathbb{U}}.$ (9)
+s.t. $\quad\mathcal{A}(\mathbf{a})\leq 0.$
+In the following, we discuss two complementary strategies to solve the optimization problem: a simplification under smooth manifold assumptions and a gradient-based optimization strategy for general cases.
+III-B Simplification under Smooth Manifold Assumption
+If the equilibrium constraint admits a smooth parametrization of joint states by the input space, i.e., $\boldsymbol{\theta}^{*}:\mathbb{U}\to\mathbb{T}$ defines a smooth immersion, then the mapping $\boldsymbol{\theta}^{*}(\boldsymbol{u})$ satisfies the implicit derivative relation:
+$\frac{\partial\boldsymbol{\theta}}{\partial\boldsymbol{u}}=-\mathbf{S}(\boldsymbol{\theta},\boldsymbol{u};\mathbf{a})^{-1}\,\mathbf{M}(\boldsymbol{\theta},\boldsymbol{u};\mathbf{a}),$ (10)
+where
+$\mathbf{S}=\frac{\partial^{2}\mathcal{E}}{\partial\boldsymbol{\theta}\partial\boldsymbol{\theta}}-\frac{\partial\boldsymbol{\tau}}{\partial\boldsymbol{\theta}},\quad\mathbf{M}=\frac{\partial^{2}\mathcal{E}}{\partial\boldsymbol{\theta}\partial\boldsymbol{u}}$ (11)
+are the generalized stiffness matrix and characterize the coupling with the input, respectively.
+The immersion property of $\boldsymbol{\theta}^{*}$ depends on the rank of $\mathbf{S}^{-1}\mathbf{M}$, the mapping is locally smooth and non-degenerate if $\mathbf{S}$ is nonsingular and $\mathbf{M}$ has full column rank.
+In this case, the feasible set $\mathbb{T}^{*}$ is diffeomorphic to the input domain $\mathbb{U}$, and the inner integral in (8) reduces to an evaluation along the immersion $\boldsymbol{\theta}^{*}(\boldsymbol{u})$.
+By applying the area formula [24], one obtains
+$\int_{\mathbb{T}^{*}}z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta};\mathbf{a}))\,\mathrm{d}\mathcal{H}^{\dim(\mathbb{T}^{*})}(\boldsymbol{\theta})=z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta}^{*};\mathbf{a}))\,\mathcal{J}_{\boldsymbol{\theta}}(\boldsymbol{u};\mathbf{a}),$
+where $\mathcal{J}_{\boldsymbol{\theta}}(\boldsymbol{u};\mathbf{a})$ is the Jacobian determinant of the immersion $\boldsymbol{\theta}^{*}(\boldsymbol{u})$, given by
+$\mathcal{J}_{\boldsymbol{\theta}}(\boldsymbol{u};\mathbf{a})=\left[\det\left(\frac{\partial\boldsymbol{\theta}}{\partial\boldsymbol{u}}^{\top}\frac{\partial\boldsymbol{\theta}}{\partial\boldsymbol{u}}\right)\right]^{1/2}\triangleq\left[\mbox{det}\left(\mathbf{G}_{\boldsymbol{\theta}}\right)\right]^{1/2}$
+where $\mathbf{G}_{\boldsymbol{\theta}}$ denotes the Gram matrix of $\frac{\partial\boldsymbol{\theta}}{\partial\boldsymbol{u}}$.
+Substituting back, the global optimization problem (8) takes the following reformulation:
+$\min_{\mathbf{a}}\mathcal{Z}(\mathbf{a})$ $=\int_{\mathbb{U}}z({}^{s}_{b}\mathbf{H}(\boldsymbol{\theta}^{*}(\boldsymbol{u});\mathbf{a}))\,\mathcal{J}_{\boldsymbol{\theta}}(\boldsymbol{u};\mathbf{a})\,\Omega_{\mathbb{U}}$ (12)
+$\mathrm{s.t.}$ $\quad\mathcal{A}(\mathbf{a})\leq 0.$
+If the structural feasibility condition ($\mathcal{A}(\mathbf{a})\leq 0$) represents bound constraints, it can be treated as part of the admissible design domain $\mathbb{A}$.
+Within this domain, the optimization reduces to an unconstrained problem, and the optimality is ensured by satisfying the following first-order necessary condition:
+$\nabla_{\mathbf{a}}\mathcal{Z}$ $=\int_{\mathbb{U}}\bigg(z\,\nabla_{\mathbf{a}}\mathcal{J}_{\boldsymbol{\theta}}+\mathcal{J}_{\boldsymbol{\theta}}\,\nabla_{\mathbf{a}}z\bigg)\,\Omega_{\mathbb{U}}$
+$=\int_{\mathbb{U}}\left[\frac{1}{2}z\mathcal{J}_{\boldsymbol{\theta}}\operatorname{tr}\left(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{\mathbf{a}}\mathbf{G}_{\boldsymbol{\theta}}\right)+\mathcal{J}_{\boldsymbol{\theta}}\nabla_{\mathbf{a}}z\right]\ \Omega_{\mathbb{U}}$
+$=\mathbf{0},$
+which, under regularity, yields the strong-form pointwise condition
+$\bigg(\frac{1}{2}z\,\operatorname{tr}(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{\mathbf{a}}\mathbf{G}_{\boldsymbol{\theta}})+\nabla_{\mathbf{a}}z\bigg)\mathcal{J}_{\boldsymbol{\theta}}=\mathbf{0},\quad\forall\boldsymbol{u}\in\mathbb{U},$ (13)
+where $\mathbf{G}_{\boldsymbol{\theta}}$ is the Gram matrix of $\tfrac{\partial\boldsymbol{\theta}}{\partial\boldsymbol{u}}$.
+Degeneracy occurs when $\mathcal{J}_{\boldsymbol{\theta}}=0$, corresponding to collapsed input directions where actuation can not produce independent generalized torques.
+If $z$ is a non-zero kinematic performance index, one can further deduce the non-trivial solution of (13) that
+$\operatorname{tr}(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{\mathbf{a}}\mathbf{G}_{\boldsymbol{\theta}})=\sum_{i}\frac{\nabla_{\mathbf{a}}\lambda_{i}(\mathbf{G}_{\boldsymbol{\theta}})}{\lambda_{i}(\mathbf{G}_{\boldsymbol{\theta}})}=-2\frac{\nabla_{\mathbf{a}}z}{z}.$ (14)
+Geometrically, this condition implies that optimizing the local performance requires structurally modulating the spectrum of the configuration space metric to counteract its distortion.
+III-C Gradient-based Optimization for General Cases
+The smooth manifold assumption may fail in practice when the generalized stiffness matrix $(\mathbf{S})$ becomes singular or ill-conditioned.
+In such cases, the implicit relation (10) is nolonger to hold, and the equilibrium mapping $\boldsymbol{\theta}^{*}$ cannot be obtained analytically.
+To address this difficulty, we reformulate the optimization as a nested energy-minimization problem that unifies both smooth and singular cases within a single computational framework.
+III-C1 Inner-level equilibrium computation
+For each input $\boldsymbol{u}$ and structural parameter $\mathbf{a}$, the equilibrium configuration is computed as the stationary point of the total potential energy:
+$\boldsymbol{\theta}^{*}(\boldsymbol{u};\mathbf{a})=\arg\min_{\boldsymbol{\theta}}\left(\mathcal{E}(\boldsymbol{\theta},\boldsymbol{u};\mathbf{a})-\int\boldsymbol{\tau}\mathrm{d}\boldsymbol{\theta}\right).$
+The minimization is solved iteratively using a trust-region scheme
+$(\mathbf{S}+\mu\mathbf{I})\Delta\boldsymbol{\theta}=\boldsymbol{\tau}(\boldsymbol{\theta})-\nabla_{\boldsymbol{\theta}}\mathcal{E},\quad\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}+\Delta\boldsymbol{\theta}.$ (15)
+The damping factor $\mu>0$ is adaptively updated based on the ratio of the actual to the predicted energy reductions.
+During the actuation, the solution for each $\boldsymbol{u}$ is warm-started from the previous equilibrium, enabling smooth continuation across potential branch points.
+III-C2 Outer-level structural update
+Once the inner equilibrium mapping is obtained, the global performance index is evaluated numerically by quadrature.
+The structural parameters are then updated following a projected gradient-descent rule:
+$\mathbf{a}_{k+1}=\mathcal{P}_{\mathbb{A}_{c}}(\mathbf{a}_{k}-\eta\nabla_{\mathbf{a}}\mathcal{Z}(\mathbf{a}_{k})),\quad\mathbb{A}_{c}=\{\mathbf{a}|\mathcal{A}(\mathbf{a})\leq 0,\mathbf{a}\in\mathbb{A}\}.$
+where $\mathcal{P}_{\mathbb{A}_{c}}(\cdot)$ denotes the projection onto the admissible design domain and $\eta>0$ is a step size ensuring monotonic decrease of $\mathcal{Z}$.
+The gradient $(\nabla_{\mathbf{a}}\mathcal{Z})$ can be approximated by finite differences or obtained via automatic differentiation through the inner minimization loop.
+Consequently, the complete computational process is summarized in Algorithm 1.
+Algorithm 1 Two-level Structural Optimization Framework
+1:Initial structural parameters $\mathbf{a}_{0}$, admissible domain $\mathbb{A}_{c}$, actuation set $\mathbb{U}$, initial configurations $\boldsymbol{\theta}_{0}(\boldsymbol{u})$, damping factor $\mu_{0}>0$.
+2:Optimized structural parameters $\mathbf{a}^{*}$.
+3:for $k=0,1,\ldots$ until convergence do
+4:  for each actuation input $\boldsymbol{u}_{j}\in\mathbb{U}$ do
+5:    Initialize $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}_{0}(\boldsymbol{u}_{j})$, $\lambda\leftarrow\lambda_{0}$.
+6:    repeat
+7:     Compute gradient $\nabla_{\boldsymbol{\theta}}\mathcal{E}(\boldsymbol{\theta},\boldsymbol{u}_{j};\mathbf{a}_{k})$
+8:     Compute the stiffness matrix $\mathbf{S}$.
+9:     Solve and update the trust-region step (15).
+10:    until convergence of $\mathcal{E}$
+11:    Store equilibrium $\boldsymbol{\theta}^{*}(\boldsymbol{u}_{j};\mathbf{a}_{k})\leftarrow\boldsymbol{\theta}$.
+12:  end for
+13:  Evaluate the global performance index $\mathcal{Z}(\mathbf{a}_{k})$.
+14:  Update structural parameters by projected gradient descent: $\mathbf{a}_{k+1}=\mathcal{P}_{\mathcal{A}_{c}}\!\big(\mathbf{a}_{k}-\eta\,\nabla_{\mathbf{a}}\mathcal{Z}(\mathbf{a}_{k})\big).$
+15:end for
+16:return $\mathbf{a}^{*}\leftarrow\mathbf{a}_{k}$.
+IV Application of the Optimization Framework to Magnetic Actuation
+To demonstrate the generality and practical relevance of the proposed optimization framework, it is applied here to MeSCRs, whose actuation arises from the interaction between internal permanent magnets and external magnetic fields.
+IV-A Magnetic Actuation
+We formulate the magnetic potential energy and its differential properties with respect to the robot configuration.
+Each embedded magnet is treated as an independent joint, with additional joints added on either side to assist the description.
 The joint indices of all $N_{m}$ micromagnets constitute an index set $\mathbb{K}\subset\mathbb{Z}$.
 Given that the robot operates under a constant temperature, the formulation aligns with the standard Helmholtz free energy without thermal variations [25].
-Additional terms with thermal effects can be found in [29].
 In this case, the magnetic potential energy of the MeSCR is given by
@@
 For readibility, we sometimes denote $\mathbf{b}(\mathbf{p}_{k})$ in $\mathbf{b}_{k_{i}}$, in which $i=0,1,\dots,N_{m}-1$.
-It is worth noting that some operational scenarios of MeSCR (e.g., interaction with human tissue) typically require low-speed motion to ensure operational safety.
-Therefore, it is generally assumed that MeSCR operates in quasi-static states, where the inertia terms can be neglected.
-The dynamic analysis of MeSCR is thus reduced to finding the equilibrium between internal/external forces and potential energy, which can be obtained by solving the following simplified Lagrangian equation:
-$\nabla_{\boldsymbol{\theta}}\mathcal{E}_{e}+\nabla_{\boldsymbol{\theta}}\mathcal{E}_{m}+\boldsymbol{\tau}(\boldsymbol{\theta})=\mathbf{0}$ (5)
-where $\boldsymbol{\tau}(\boldsymbol{\theta})$ is the internal/external applied torques that can be considered (e.g., gravity torques).
-We further provide the analytical expressions for the gradients and Hessians of the potential energies, which can improve the accuracy and efficiency of calculations.
 The gradient and Hessian of $\mathcal{E}_{e}$ are straightforward, whereas we focus on computing that of $\mathcal{E}_{m}$.
-$\nabla_{\boldsymbol{\theta}_{i}}\mathcal{E}_{m}=$ $-\sum_{k\in\mathbb{K},k\geq i}\left(\left[\mathbf{m}_{k}\right]_{\times}^{\top}\mathbf{R}_{0}^{i}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}\right)^{\top}\mathbf{b}(\mathbf{p}_{k})$ (6)
-$-\sum_{k\in\mathbb{K},k\geq i}\left(\frac{\partial\mathbf{b}(\mathbf{p}_{k})}{\partial\mathbf{p}_{k}}\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}\right)^{\top}\mathbf{m}_{k}$
-$\triangleq$ $-\underbrace{\mathbf{M}_{i}(\boldsymbol{\theta})\mathbf{b}}_{\begin{subarray}{c}\text{magnetic}\\ \text{torques}\end{subarray}}-\sum_{k\in\mathbb{K},k\geq i}\left(\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}\right)^{\top}\underbrace{\left(\nabla_{\mathbf{p}_{k}}\mathbf{b}\right)\mathbf{m}_{k}}_{\begin{subarray}{c}\text{magnetic}\\ \text{forces}\end{subarray}}$
+$\nabla_{\boldsymbol{\theta}_{i}}\mathcal{E}_{m}=$ $-\sum_{k\in\mathbb{K},k\geq i}\left(\left[\mathbf{m}_{k}\right]_{\times}^{\top}\mathbf{R}_{0}^{i}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}+\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}(\mathbf{m}_{k}^{\top}\nabla_{\mathbf{p}_{i}})\right)^{\top}\mathbf{b}_{k}$
+$\triangleq$ $\mathbf{M}_{i}(\boldsymbol{\theta})\mathbf{b}$
 where $\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}$ represents the right Jacobian of $\mathrm{SO}(3)$ manifold associated with $\boldsymbol{\theta}_{i}$ [26].
@@
 For precise calculations, $\frac{\partial\mathbf{p}_{k}}{\partial\boldsymbol{\theta}_{i}}$ can be derived from (4).
-However, the magnetic force is usually negligible in comparison to the magnetic torques, which implies the following explicit representation of (5) as
+As a result, the equilibrium condition (6) has the following explicit representation:
 $\boldsymbol{\Lambda}(\boldsymbol{\theta}-\bar{\boldsymbol{\theta}})+\boldsymbol{\tau}(\boldsymbol{\theta})=\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$ (16)
@@
 By developing (16) and performing some rearrangements, the joint space compliance relation between the variation of the actuating magnetic field $(\delta\mathbf{b})$ and the configuration variation $(\delta\boldsymbol{\theta})$ is obtained:
-$\underbrace{\left(\boldsymbol{\Lambda}+\frac{\partial^{2}\mathcal{E}_{m}}{\partial\boldsymbol{\theta}\partial\boldsymbol{\theta}^{\top}}+\frac{\partial\boldsymbol{\tau}}{\partial\boldsymbol{\theta}}\right)}_{\triangleq\mathbf{S}}\delta\boldsymbol{\theta}=\mathbf{M}(\boldsymbol{\theta})\delta\mathbf{b}$
-where the Hessian of $\mathcal{E}_{m}$ (denoted in $\mathbf{S}_{m}$) is given in the supplementary material and $\mathbf{S}\in\mathbb{R}^{3N\times 3N}$ denotes the Hessian matrix of stiffness.
-Hence, the joint-space compliance relation reads
-$\delta\boldsymbol{\theta}=\mathbf{S}^{-1}\mathbf{M}(\boldsymbol{\theta})\delta\mathbf{b}.$ (8)
-Premultiplying (8) by the space Jacobian (2) yields the Cartesian compliance mapping from magnetic field perturbations to distal pose variations:
-$\delta\mathbf{h}=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{S}^{-1}\mathbf{M}(\boldsymbol{\theta})\delta\mathbf{b}\triangleq\mathbf{J}_{\mathbf{b}}\delta\mathbf{b}$ (9)
-where $\mathbf{J}_{\mathbf{b}}\in\mathbb{R}^{6\times 3N_{m}}$ is referred to as the actuation Jacobian. We close this subsection by introducing two useful propositions:
+$\left(\boldsymbol{\Lambda}+\frac{\partial^{2}\mathcal{E}_{m}}{\partial\boldsymbol{\theta}\partial\boldsymbol{\theta}^{\top}}+\frac{\partial\boldsymbol{\tau}}{\partial\boldsymbol{\theta}}\right)\delta\boldsymbol{\theta}\triangleq\mathbf{S}\delta\boldsymbol{\theta}=\mathbf{M}(\boldsymbol{\theta})\delta\mathbf{b}$
+where $\mathbf{S}\in\mathbb{R}^{3N\times 3N}$ is known as the Hessian matrix of stiffness.
+Observed that, the magnetic actuation of the MeSCR follows the canonical equilibrium structure of (11), where variations in the magnetic field $\mathbf{b}$ correspond to the control inputs $\boldsymbol{u}$ that drive configuration changes.
+IV-B Properties of Magnetic Actuation under Uniform Fields
+We further analyze the properties of the actuation mapping in a uniform magnetic field, where the field can be expressed as $\mathbf{b}=\mathbf{U}_{u}\mathbf{b}_{u}$ with $\mathbf{U}_{u}=\mathbf{1}_{N_{m}}\otimes\mathbf{I}_{3}$ and $\mathbf{b}_{u}\in\mathbb{R}^{3}$.
+We begin by establishing two fundamental properties of the magnetic actuation mapping:
 Proposition 1.
@@
 Proof.
-The proofs are deferred to Appendix A. ∎
-II-C Analysis and Validation of Equilibrium Solution
-[…]
-TABLE I: Geometric and Material properties of the MeSCR#1
-[…]
-[…]
-[…]
-[…]
-$I$ (Area moment of inertia) $0.7854$ $[\mbox{mm}^{4}]$
-[…]
-[…]
-$M$ (Modulus of the magnetic dipole moment) $10^{-2}$ $[\mbox{A}\cdot\mbox{m}^{2}]$
-We first analyze the equilibrium solution.
-Assuming the robot to be naturally straight $(\bar{\boldsymbol{\theta}}=\mathbf{0})$, and $\boldsymbol{\tau}(\boldsymbol{\theta})$ is ruled out.
-Then, reformulate (7) derives the following fixed-point equation:
-$\boldsymbol{\theta}=\boldsymbol{\Lambda}^{-1}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}.$ (10)
-Observe that the right-hand side of the equation represents a continuous mapping with respect to $\boldsymbol{\theta}$. We present the following theorem regarding the uniqueness of the solution.
+The proof is deferred to supplementary materials. ∎
+Assuming the robot to be naturally straight $(\bar{\boldsymbol{\theta}}=\mathbf{0})$, and $\boldsymbol{\tau}(\boldsymbol{\theta})$ is ruled out, (16) further simplifies to the following fixed-point equation:
+$\boldsymbol{\theta}=\boldsymbol{\Lambda}^{-1}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$ (17)
+Under this formulation, the existence and uniqueness of the equilibrium solution can be guaranteed by ensuring that the mapping is a contraction in the configuration space:
 Theorem 1.
@@
 The inequality provides a sufficient condition to guarantee the uniqueness of a globally attractive equilibrium point for any initial condition.
-Notably, the upper bound is solely determined by the material properties: $E$, $I$, $L$, $M_{k}$, and is independent of the joint number $N$ (note that $N\bar{\ell}_{i}\approx L)$.
+Notably, the upper bound is solely determined by the material properties and is independent of the joint number $N$ (note that $N\bar{\ell}_{i}\approx L$).
 The limitation on the external magnetic field strength can be explained by the critical buckling instability of the magneto-elastic material [27].
-We found it intriguing that when each joint is magnetized, the resulting upper bound exhibits a striking structural resemblance to that derived in our previous work [32] on magnetic continua, differing only slightly in the coefficients.
-This similarity highlights the underlying consistency in the physical principles governing continuum and discreteness.
+We found it intriguing that when each joint is magnetized, the resulting upper bound exhibits a striking structural resemblance to that derived in previous work [28] on magnetic continua, differing only slightly in the coefficients.
 Furthermore, Theorem 1 also ensures $\mathbf{S}$ to be invertible.
 By applying Weyl’s inequality, one can easily verify that $\lambda_{\min}(\mathbf{S})\geq\lambda_{\min}(\boldsymbol{\Lambda})-\mathcal{L}>0$.
-Although the upper bound remains a conservative estimate, once the external magnetic field strength exceeds this upper limit, (10) may admit multiple equilibria pertaining to different initial conditions [21].
-Therefore, selecting appropriate initial values for the iteration becomes crucial to ensuring convergence and avoiding undesirable behaviors.
-The solution trajectory can be described by the constructed dynamic system $\frac{\mathrm{d}\boldsymbol{\theta}}{\mathrm{d}t}=-\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}}$ with initial condition $\boldsymbol{\theta}(0)$, which can also be regarded as a gradient flow that minimizes the energy function $\mathcal{E}(\boldsymbol{\theta})$.
-We propose the following theorem and lemmas, which guarantee that the solution is bounded and convergent by choosing an appropriate initial condition.
-Lemma 1.
-For any initial condition $\boldsymbol{\theta}(0)$, the solution trajectory $\boldsymbol{\theta}(t)$ is bounded that
-$\|\boldsymbol{\theta}(t)\|\leq\max\left\{\|\boldsymbol{\theta}(0)\|,\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\right\}.$
-Proof.
-Consider the Lyapunov function $V(\boldsymbol{\theta})=\frac{1}{2}\|\boldsymbol{\theta}\|^{2}$, the derivative of $V(\boldsymbol{\theta})$ with respect to time is $\dot{V}(\boldsymbol{\theta})=\boldsymbol{\theta}^{\top}(-\boldsymbol{\Lambda}\boldsymbol{\theta}+\mathbf{M}(\boldsymbol{\theta})\mathbf{b})=-\boldsymbol{\theta}^{\top}\boldsymbol{\Lambda}\boldsymbol{\theta}+\boldsymbol{\theta}^{\top}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$.
-Using Proposition 1, the upper bound of $\dot{V}(\boldsymbol{\theta})$ is derived as $\dot{V}(\boldsymbol{\theta})\leq-\lambda_{\min}(\boldsymbol{\Lambda})\|\boldsymbol{\theta}\|^{2}+\mathcal{M}\|\boldsymbol{\theta}\|.$ Let $W(t)=\sqrt{2V(\boldsymbol{\theta})}=\|\boldsymbol{\theta}\|$, one can find out that: $\dot{W}(t)=\frac{\dot{V}(\boldsymbol{\theta})}{W(t)}\leq-\lambda_{\min}W(t)+\mathcal{M}.$ Applying Grönwall’s inequality, we obtain $W(t)\leq W(0)e^{-\lambda_{\min}(\boldsymbol{\Lambda})t}+\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\left(1-e^{-\lambda_{\min}\left(\boldsymbol{\Lambda}\right)t}\right).$ Observed that as $t\to\infty$, the exponential term vanishes, leading to $\lim_{t\to\infty}W(t)\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}.$ This completes the proof, establishing that $\|\boldsymbol{\theta}(t)\|$ remains $\|\boldsymbol{\theta}(t)\|\leq\max\left\{\|\boldsymbol{\theta}(0)\|,\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\right\}.$
-Lemma 2.
-Let $\boldsymbol{\theta}^{*}$ be any equilibrium point of (10).
-Using Proposition 1, we have $\|\boldsymbol{\theta}^{*}\|=\|\boldsymbol{\Lambda}^{-1}\mathbf{M}(\boldsymbol{\theta}^{*})\mathbf{b}\|\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}$.
-Theorem 2.
-The solution trajectory $(\boldsymbol{\theta}(t))$ will always converge to the largest invariant subset of the equilibrium point set if the initial condition satisfies $\|\boldsymbol{\theta}(0)\|\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}$.
-Proof.
-Choose the Lyapunov function $V(\boldsymbol{\theta})=\mathcal{E}(\boldsymbol{\theta})$, whose derivative along the solution trajectory satisfies $\dot{V}(\boldsymbol{\theta})=-\|\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}}\|^{2}\leq 0$, with equality holding if and only if $\boldsymbol{\theta}$ is an equilibrium point.
-Consider the positively invariant set $\Omega=\left\{\boldsymbol{\theta}\in\mathbb{R}^{3N}\mid\|\boldsymbol{\theta}\|\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\right\}$ (Lemma 1), which contains all equilibrium points (Lemma 2).
-By Brouwer’s fixed-point theorem, there exists at least one equilibrium point within $\Omega$, that is, $\dot{V}(\boldsymbol{\theta})\leq 0$ holds for all $\boldsymbol{\theta}\in\Omega$.
-According to LaSalle’s invariance principle, every trajectory within $\Omega$ converges to the largest invariant subset where $\dot{V}(\boldsymbol{\theta})=0$.
-∎
-When the external magnetic field changes, requiring the recalculation of the equilibrium point, Theorem 2 provides a natural choice for the initial condition as $\boldsymbol{\theta}(0)=\mathbf{0}$ because $\forall\mathbf{b}\in\mathbb{B},\mathcal{M}\geq 0$.
-One can also choose the nearby solution as an initial guess if $\mathbf{b}$ continuously increases.
-For small perturbations on $\mathbf{b}$, the corresponding $\delta\boldsymbol{\theta}$ can be calculated by (8).
-Moreover, from Proposition 1 and the fact that $\boldsymbol{\theta}=\mathbf{0}$ is the unique equilibrium solution when $\mathbf{b}=\mathbf{0}$, we deduce that $\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}}$ is Lipschitz continuous with respect to $\mathbf{b}$, and the equilibrium points depends continuously on it.
-[…]
-Notably, using fewer joints does not introduce significant truncation errors and offers a computational speed advantage over BVP-solving models.
-Fig.
-3e and Fig.
-3f show the root-mean-square error (RMSE) of the centerline position and the distal pose error norm as the joint count increases from $3$ to $100$, benchmarked against $200$ joints.
-Results demonstrate that RMSE decreases proportionally to $1/N^{2}$, with an error bound established in our previous work [34].
-Using $7$ joints achieves an RMSE below the acceptable $10^{-4}$ threshold.
-Additionally, the distal pose error norm declines rapidly as $N$ grows, confirming that fewer joints can accurately compute the distal pose under external magnetic fields.
-This analysis leads to a key insight: the performance index of the actuation Jacobian is largely insensitive to $N$ and can be understood as an inherent property of the robot in a specific configuration.
-III Enhancing Kinematic Performances
-This section focuses on the structural optimization of MeSCRs to enhance kinematic performances, including manipulability and dexterity.
-Prior to the optimization, we perform some kinematic analysis of the actuation Jacobian, uncovering how different magnet configuration shape the robot’s controllable DoF achievable workspace.
-III-A Kinematics Analysis
-We define the controllable degrees of freedom (DoF) as the independently actuated direction of the task-space twist, which corresponds to the row rank of the actuation Jacobian ($\mathbf{J}_{\mathbf{b}}$).
+We subsequently define the controllable degrees of freedom (DoF) as the independently actuated direction of the task-space twist, which corresponds to the row rank of the actuation Jacobian $(\mathbf{J}_{\mathbf{b}}=\mathbf{J}_{\boldsymbol{\theta}}\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}})$.
 The following theorem reveals the relation between it and the configuration of embedded magnets:
@@
 Proof.
-Recall that the actuation Jacobian $\mathbf{J}_{\mathbf{b}}=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{S}^{-1}\mathbf{M}(\boldsymbol{\theta})$, the post-multiplied matrix $(\mathbf{M}(\boldsymbol{\theta}))$ can be rewritten as
-$\mathbf{M}(\boldsymbol{\theta})$ $=\mathbf{D}_{k_{N_{m}-1}}^{\top}\left(\begin{bmatrix}\begin{matrix}\mathbf{1}_{k_{0}}\\ \mathbf{0}_{N-k_{0}}\end{matrix}\cdots\begin{matrix}\mathbf{1}_{k_{N_{m}-1}}\\ \mathbf{0}_{N-k_{N_{m}-1}}\end{matrix}\end{bmatrix}\otimes\mathbf{I}_{3\times 3}\right)$ (12)
-×$\times\operatorname{blkdiag}\left\{\left[\mathbf{m}_{k_{0}}\right]_{\times},\dots,\left[\mathbf{m}_{k_{N_{m}-1}}\right]_{\times}\right\}$
-in which the block diagonal matrix $\mathbf{D}_{k_{N_{m}-1}}$ is provided in (24) and $\mathbf{1}_{k}$ denotes the all-ones vector of $k$ dimensions.
-To determine the rank of $\mathbf{M}(\boldsymbol{\theta})$, one starts by noting that
-$\operatorname{rk}\left(\begin{bmatrix}\begin{matrix}\mathbf{1}_{k_{0}}\\ \mathbf{0}_{N-k_{0}}\end{matrix}\cdots\begin{matrix}\mathbf{1}_{k_{N_{m}-1}}\\ \mathbf{0}_{N-k_{N_{m}-1}}\end{matrix}\end{bmatrix}\otimes\mathbf{I}_{3}\right)=3N_{m}.$
-Applying Sylvester’s rk inequality, we obtain:
-$\operatorname{rk}\left(\mathbf{M}(\boldsymbol{\theta})\right)$ $\leq\min\!\left\{\sum_{i=0}^{N_{m}-1}\!\operatorname{rk}\!\left(\mathbf{R}_{0}^{k_{i}}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k_{i}}}\right),\sum_{i=0}^{N_{m}-1}\!\operatorname{rk}\!\left(\left[\mathbf{m}_{k_{i}}\right]_{\times}\right)\right\}$ (13)
-$=2N_{m},$
-and
-$\operatorname{rk}\!\left(\mathbf{M}(\boldsymbol{\theta})\right)$ $\geq\sum_{i=0}^{N_{m}-1}\!\left(\operatorname{rk}\!\left(\mathbf{R}_{0}^{k_{i}}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k_{i}}}\right)+\!\operatorname{rk}\!\left(\left[\mathbf{m}_{k_{i}}\right]_{\times}\right)\right)-3N_{m}$ (14)
-$=2N_{m}.$
-Here we used the facts that $\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k_{i}}}$ and $\mathbf{R}_{0}^{k_{i}}$ are both invertible and the rank of a skew-symmetric matrix remains $2$.
-Combined (13) with (14) leads to the following result:
-$\operatorname{rk}\left(\mathbf{M}(\boldsymbol{\theta})\right)=2N_{m}.$
-Consequently, an important conclusion is derived as $\operatorname{rk}\left(\mathbf{J}_{\mathbf{b}}\right)\leq\min\left\{6,\operatorname{rk}\left(\mathbf{M}(\boldsymbol{\theta})\right)\right\}=\min\left\{6,2N_{m}\right\}$.
-∎
-We proceed with a brief discussion on MeSCR’s structural configuration with different $N_{m}$:
-Figure 4: Deformation of the MeSCR with different numbers of embedded magnets $(N_{m}=1,2,3,4)$, corresponding to subfigures a.), b.), c.), and d.).
-As $N_{m}$ increases, the MeSCR is characterized as underactuated for $N_{m}=1,2$, fully-actuated for $N_{m}=3$, and redundant for $N_{m}=4$.
-  • $N_{m}=1$ [2, 3, 4] (see Fig.
-4a.): The simplest yet most used configuration.
-The single magnet is typically embedded at the distal end to maximize torque transmission.
-From Theorem 3, we show that at most $2$ DoF of the task-space twist is controllable, which is consistent with the conclusion in [18].
-  • $N_{m}=2$[35, 20, 19, 36] (see Fig.
-4b.): Multimodality properties emerged since the orientation of embedded magnets is designable.
-However, at most $4$ DoF are controllable, which still characterizes the system as underactuated.
-  • $N_{m}=3$ [37, 13, 38] (see Fig. 4c.): The structural configuration of the embedded magnets is also optimizable. The maximum controllable DoF is $6$, indicating that the MeSCR is fully-actuated.
-  • $N_{m}\geq 4$[39, 40, 41] (see Fig.
-4d.): Adopted by many commerical products.
-The column rank of $\mathbf{J}_{\mathbf{b}}$ exceeds its row rank, indicating that the MeSCR is a redundant system and the number of redundant DoF is $62N_{m}-$ 6.
-The maximum controllable DoF will also be limited by the actuating field.
-As $N_{m}$ increases, achieving fully-actuated or redundant control requires each embedded magnet to be paired with an independent actuating field.
-However, generating independently controllable actuating fields at arbitrary positions within the workspace is highly challenging.
-While some advanced equipment may be capable of this [42], it is not feasible for most conventional systems.
-For example, when the actuating field is spatially uniform within the workspace, which can be expressed as $\mathbf{b}=\mathbf{U}_{u}\mathbf{b}_{u}$ with $\mathbf{U}_{u}=\mathbf{1}_{N_{m}}\otimes\mathbf{I}_{3}$ and $\mathbf{b}_{u}\in\mathbb{R}^{3}$.
-In this case, the maximum controllable DoF will not exceed $3$.
-To reduce the number of design variables, we assume all embedded magnets have uniform material and geometric properties, which also facilitates the manufacturing process.
-Consequently, the orientation and position of the embedded magnets become the primary variables to be optimized.
-In existing commercial products, embedded magnets are generally placed along the catheter’s axial direction.
-For $N_{m}$ magnets, this arrangement leads to $2^{N_{m}}$ possible permutations.
-However, if the actuating magnetic field of each magnet is independently controllable, the directions of the magnetic moments become redundant.
-This redundancy arises because of the relationship: [±$[\pm\mathbf{m}_{k}]_{\times}\mathbf{b}_{k}=[\mathbf{m}_{k}]_{\times}(\pm\mathbf{b}_{k})$, indicating that the effect of flipping a magnetic moment’s direction can always be compensated by reversing the corresponding magnetic field direction.
-Based on this reasoning, we assume that the magnetic field can only undergo an overall direction reversal.
-For systems driven by electromagnets, this can be easily achieved by altering the current direction.
-Aligning the magnetic moments along the axial direction significantly reduces the optimization search space but imposes a limitation on the motion of the MeSCR.
-Specifically, this arrangement restricts the MeSCR to material-twist-free due to the decoupling of twisting and bending, as ensured by the isotropic and axially symmetric properties of its soft matrix.
+It is deferred to the supplementary materials. ∎
+The controllable DoF is ultimately constrained by the number and independence of the actuating fields.
+As $N_{m}$ increases, achieving fully actuated or redundant control would require an equal number of independently controllable fields, which is generally infeasible in practice.
+Under the spatially uniform condition on fields, the maximum controllable DoF does not exceed three.
+For simplicity and manufacturability, all embedded magnets are assumed to share identical material and geometric properties.
+In most commercial designs, embedded magnets are placed along the catheter’s axial direction.
+For $N_{m}$ magnets, this yields $2^{N_{m}}$ orientation permutations.
+However, because $\pm\mathbf{m}_{k}]_{\times}\mathbf{b}_{k}=[\mathbf{m}_{k}]_{\times}(\pm\mathbf{b}_{k})$, flipping a magnet’s polarity can be equivalently compensated by reversing the field direction.
+Thus, axial alignment effectively reduces the optimization space while restricting the MeSCR to a twist-free deformation mode due to the isotropic and axially symmetric nature of its soft matrix.
 Theorem 3 (Material-twist-free).
-The MeSCR is material-twist-free ($\boldsymbol{\theta}_{i}^{\top}\bar{\mathbf{t}}_{i}=0$) for any actuating magnetic field, provided that the embedded magnets’ magnetic moments are aligned axially, i.e., $\bar{\mathbf{m}}_{k}=\pm M_{k}\bar{\mathbf{t}}$, and the magnetic field strength satisfies:
-$\max_{k\in\mathbb{K}}B_{k}\leq\frac{\sqrt{6}\lambda_{\min}(\boldsymbol{\Lambda})}{\sum_{k\in\mathbb{K}}M_{k}}.$ (15)
+The MeSCR is material-twist-free ($\boldsymbol{\theta}_{i}^{\top}\bar{\mathbf{t}}_{i}=0$) for any actuating magnetic field, provided that the embedded magnets’ magnetic moments are aligned axially, i.e., $\bar{\mathbf{m}}_{k}=\pm M_{k}\bar{\mathbf{t}}$, and the magnetic field strength satisfies: $\max_{k\in\mathbb{K}}B_{k}\leq\sqrt{6}\lambda_{\min}(\boldsymbol{\Lambda})/\sum_{k\in\mathbb{K}}M_{k}.$
 Proof.
-The proof is detailed in the supplementary material. ∎
+See supplementary material for details. ∎
 Noticing that $\frac{\sqrt{6}\lambda_{\min}(\boldsymbol{\Lambda})}{\sum_{k\in\mathbb{K}}M_{k}}\propto N^{2}$, as $N\to\infty$, Theorem 3 holds for any actuating magnetic field.
-This upper bound characterizes the ability of the discrete model with $N$ joints to approximate the continuum, as discussed in Section II-C.
+This upper bound characterizes the ability of the discrete model with $N$ joints to approximate the continuum.
 We found an alternative form of the equilibrium equation (17) from the underlying proof:
 $\boldsymbol{\theta}_{i}^{*}=\frac{\sum_{k\geq i}\mathbf{v}_{k}}{\lambda_{\max}(\boldsymbol{\Lambda}_{i})}=\frac{\sum_{k\geq i}[\mathbf{R}_{0}^{k^{\top}}(\boldsymbol{\theta}^{*})\mathbf{b}_{k}]_{\times}^{\top}\bar{\mathbf{m}}_{k}}{\lambda_{\max}(\boldsymbol{\Lambda}_{i})}.$ (19)
-For any initial condition satisfying the assumptions of Theorem 2, the associated equilibrium state is guaranteed to exhibit zero torsional components.
 The configuration space (1) thus can be degenerated to a simplified model as presented in [29].
@@
 Corollary 1.
-Define the plane $\Pi=\operatorname{span}\{\bar{\mathbf{t}},\mathbf{v}\}\subset\mathbb{R}^{3}$, where $\mathbf{v}\in\mathbb{R}^{3}$ is linearly independent of $\bar{\mathbf{t}}$.
+Define the plane $\Pi=\operatorname{span}\{\bar{\mathbf{t}},\mathbf{s}\}\subset\mathbb{R}^{3}$, where $\mathbf{s}\in\mathbb{R}^{3}$ is linearly independent of $\bar{\mathbf{t}}$.
 Suppose $\forall k\in\mathbb{K}$, $\bar{\mathbf{m}}_{k}=\pm M_{k}\bar{\mathbf{t}}$ and $\mathbf{b}_{k}\in\Pi$.
@@
 Proof.
-[…]
+The proof is constructive.
+Let $\boldsymbol{\theta}_{i}^{*}=\theta_{i}^{*}\mathbf{n}$, where $\mathbf{n}=\frac{\bar{\mathbf{t}}\times\mathbf{s}}{\|\bar{\mathbf{t}}\times\mathbf{s}\|}$ is the unit vector orthogonal to $\Pi$.
+Since $\mathbf{b}_{k}\in\Pi$, we notice that $\mathbf{R}_{0}^{k^{\top}}(\boldsymbol{\theta}^{*})\mathbf{b}_{k}=\mathbf{R}(-\sum_{i}\theta_{i}^{*}\mathbf{n})\mathbf{b}_{k}\in\Pi$.
+Then $\forall k\in\mathbb{K}$, $\bar{\mathbf{m}}_{k}]_{\times}\mathbf{R}_{0}^{k^{\top}}\mathbf{b}_{k}\!\in\operatorname{span}\{\mathbf{n}\}$.
+Hence $\boldsymbol{\theta}_{i}^{*}=\theta_{i}^{*}\mathbf{n}$ is a solution to (19).
+The corresponding tangent vectors $\mathbf{t}_{i}=\mathbf{R}_{0}^{i}\bar{\mathbf{t}}_{i}\in\Pi$, thus the centerline lies entierly within the plane $\Pi$.
+∎
 Specially, if the initial value is chosen as $\boldsymbol{\theta}(0)=\mathbf{0}$, then the gradient-based iteration evolves within the subspace $\operatorname{span}\{\mathbf{n}\}$.
-By Theorem 2, the optimization trajectory will converge to a planar equilibrium configuration satisfying $\boldsymbol{\theta}_{i}=\theta_{i}\mathbf{n}$.
+The optimization trajectory converges to a planar equilibrium configuration satisfying $\boldsymbol{\theta}_{i}=\theta_{i}\mathbf{n}$.
 We can further derive a closed-form expression under the small-angle premise, wherein the cumulative rotation matrix admits the first-order expansion: $\mathbf{R}\left(\sum_{i=0}^{k}\theta_{i}\mathbf{n}\right)\approx\mathbf{I}_{3}+\sum_{i=0}^{k}\theta_{i}[\mathbf{n}]_{\times}$.
@@
 The optimization problem $\min_{\boldsymbol{\theta}}\mathcal{E}$ thus has the explicit solution
-$\theta_{i}=\frac{1}{\Lambda_{i}}\sum_{k\in\mathbb{K},k\geq i}\bar{\mathbf{m}}_{k}^{\top}[\mathbf{b}_{k}]_{\times}\mathbf{n}.$ (17)
+$\theta_{i}=\frac{1}{\Lambda_{i}}\sum_{k\in\mathbb{K},k\geq i}\bar{\mathbf{m}}_{k}^{\top}[\mathbf{n}]_{\times}^{\top}\mathbf{b}_{k}.$ (20)
 It can be verified that this also corresponds to the general solution of (19) under the small-angle premise.
-III-B Optimization Methodology
-Figure 5: Comparison of the half-planar normalized workspace of MeSCRs with equidistant placement of three embedded magnets.
-The position and orientation of embedded magnets determine the magnetic response deformation of MeSCRs under the same magnetic field.
+IV-C Analytical Optimization under Weak Uniform Fields
+If the external magnetic field remains unchanged, the deformation of MeSCRs is determined by the placement (position) and orientation of embedded magnets.
 For instance, embedded magnets with variable positions have been employed to expand the MeSCR’s workspace [30].
-In the PRB model, the length of embedded magnets can be defined by the Voronoi length as $L_{k}\triangleq\sum_{i\leq k}\ell_{i},\forall k\in\mathbb{K}$.
+In the PRB model, the length of embedded magnets is defined by the Voronoi length as $L_{k}\triangleq\sum_{i\leq k}\ell_{i},\forall k\in\mathbb{K}$.
 From (17), it follows that the flexible segment beyond the last embedded magnet is unactuated.
-Consequently, it does not affect the workspace volume of the MeSCR, nor any performance metric defined using a left-invariant Riemannian metric.
-Therefore, one of the magnets is fixed at the distal end of the continuum ($L_{k_{N_{m}-1}}=L$), leaving $N_{m}-1$ decision variables.
-Additionally, let the flexible segments between adjacent magnets be uniformly distributed and have appropriate spacing to mitigate magnetic coupling effects.
-Based on the previous assumption regarding the overall reversal of the magnetic field direction, any configuration of axially placed magnets has a corresponding symmetrical configuration, resulting in a total of $2^{N_{m}-1}$ possible combinations.
-Expanding the workspace volume is a key objective in the optimal design of MeSCRs [43].
-However, comparing workspace volumes across MeSCRs with different embedded magnet orientations is of limited significance, unless the focus is explicitly on spatial coverage by the distal end.
-As shown in Fig.
-5, we compared the normalized half-workspace areas of the four MeSCRs under equidistant configurations.
-The results indicate that ② achieves the largest normalized workspace area, whereas ③ has the smallest.
-Nevertheless, despite its smaller workspace area, ① may still reach regions that are inaccessible to ②.
-Therefore, our focus shifts to the global indices of MeSCRs with fixed magnet orientations within their respective workspaces.
-Recall that ${}_{s}^{b}\mathbf{H}(\boldsymbol{\theta}):\mathbb{T}\to\mathrm{SE}(3)$ is the forward kinematic map of the MeSCR.
-Let $z({}_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k}):\mathbb{T}\to\mathbb{R}$ be an integrable scalar functional defined on $\mathbb{T}$, representing a local performance density of ${}_{s}^{b}\mathbf{H}(\boldsymbol{\theta})$ with left-invariant Riemannian metric.
-This local measure can be extended to a global measure $\mathcal{Z}(L_{k})$ by integrating $z$ over $\mathbb{T}$, yielding the following optimization problem on $L_{k}$:
-$\min_{L_{k}}$ $\ \mathcal{Z}(L_{k})\triangleq\int_{\mathbb{T}}z({}_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\ \Omega_{\mathbb{T}}$ (18)
-[…]
-where $\Omega_{\mathbb{T}}$ represents the Riemannian volume form on $\mathbb{T}$ induced from the Euclidean metric and $\mathcal{C}(L_{k})$ denotes linear constraints on $L_{k}$.
-The equilibrium constraint restricts feasible configurations to the solution set $\mathbb{T}^{*}\triangleq\{\boldsymbol{\theta}\in\mathbb{T}|\nabla_{\boldsymbol{\theta}}\mathcal{E}(\boldsymbol{\theta},\mathbf{b})=\mathbf{0}\}$.
-If the actuation manifold $\mathbb{B}$ is arbitrary, the set $\mathbb{T}^{*}$ may not admit an explicit parametrization.
-Nevertheless, under mild regularity conditions on $\mathcal{E}$, and by Sard’s theorem, the solution set $\mathbb{T}^{*}\subset\mathbb{T}$ is countably $d$-rectifiable, with $d=3N-\operatorname{rk}(\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}})$.
-Hence, it admits a well-defined $\mathcal{H}^{d}$-measure and supports geometric integration.
-Consequently, the constrained integral over $\mathbb{T}$ may be rewritten as an iterated integral over $\mathbb{B}$ and the corresponding solution manifold:
-$\mathcal{Z}(L_{k})=\int_{\mathbb{B}}\left(\int_{\mathbb{T}^{*}}z(_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\mathrm{d}\mathcal{H}^{d}(\boldsymbol{\theta})\right)\Omega_{\mathbb{B}}.$
-Now, suppose the external magnetic field $\mathbf{b}$ is spatially uniform with $\mathbf{b}=\mathbf{U}_{u}\mathbf{b}_{u}$ and satisfies the condition in Theorem 1.
-The manifold $\mathbb{B}$ is now an affine subspace diffeomorphic to $\mathbb{R}^{3}$.
-By Theorem 1, the stiffness matrix $\mathbf{S}$ is nonsingular in a neighborhood of each $\mathbf{b}_{u}\in\mathbb{B}$.
-By Theorem 3, the Jacobian $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}}=\mathbf{S}^{-1}\mathbf{M}$ satisfies $\operatorname{rk}(\mathbf{S}^{-1}\mathbf{M})=2N_{m}$, which ensures that $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}=\mathbf{S}^{-1}\mathbf{M}\mathbf{U_{u}}$ is of full rank provided $N_{m}\geq 2$.
-In this regime, the equilibrium constraint gives rise to a smooth immersion $\boldsymbol{\theta}^{*}:\mathbb{B}\to\mathbb{T}$.
-The inner integral reduces to an evaluation at $\boldsymbol{\theta}^{*}(\mathbf{b})$, and using the area formula [44] yields:
-$\int_{\mathbb{T}^{*}}z(_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\mathrm{d}\mathcal{H}^{d}(\boldsymbol{\theta})=z(_{s}^{b}\mathbf{H}(\boldsymbol{\theta});L_{k})\mathcal{J}_{\boldsymbol{\theta}}(\mathbf{b}),$
-where $\mathcal{J}_{\boldsymbol{\theta}}(\mathbf{b})$ denotes the Jacobian of the immersion $\boldsymbol{\theta}^{*}(\mathbf{b})$ and is given by
-$\mathcal{J}_{\boldsymbol{\theta}}=\left[\det\left(\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}^{\top}\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}\right)\ \right]^{\frac{1}{2}}\triangleq\left[\mbox{det}\left(\mathbf{G}_{\boldsymbol{\theta}}\right)\right]^{\frac{1}{2}}.$
-where $\mathbf{G}_{\boldsymbol{\theta}}$ denotes the Gram matrix of $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}$.
-Substituting back, the global optimization problem (18) takes the following unconstrained reformulation:
-$\min_{L_{k}}$ $\ \mathcal{Z}(L_{k})=\int_{\mathbb{B}}z({}_{s}^{b}\mathbf{H}(\boldsymbol{\theta}^{*}(\mathbf{b}));L_{k})\mathcal{J}_{\boldsymbol{\theta}}\ \Omega_{\mathbb{B}}.$ (19)
-The optimality is ensured by satisfying the following first-order necessary condition:
-$\nabla_{L_{k}}\mathcal{Z}$ $=\int_{\mathbb{B}}\left(z\nabla_{L_{k}}\mathcal{J}_{\boldsymbol{\theta}}+\mathcal{J}_{\boldsymbol{\theta}}\nabla_{L_{k}}z\right)\ \Omega_{\mathbb{B}}$
-$=\int_{\mathbb{B}}\left[\frac{1}{2}z\mathcal{J}_{\boldsymbol{\theta}}\operatorname{tr}\left(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{L_{k}}\mathbf{G}_{\boldsymbol{\theta}}\right)+\mathcal{J}_{\boldsymbol{\theta}}\nabla_{L_{k}}z\right]\ \Omega_{\mathbb{B}}$
-$=0,\ \forall k\in\mathbb{K}/\{N_{m}-1\}.$
-Under regularity conditions, this implies the strong-form pointwise optimality condition:
-$\left(\frac{1}{2}z\operatorname{tr}\left(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{L_{k}}\mathbf{G}_{\boldsymbol{\theta}}\right)+\nabla_{L_{k}}z\right)\mathcal{J}_{\boldsymbol{\theta}}=0,\ \forall\mathbf{b}_{u}\in\mathbb{B}.$ (20)
-Observed that the condition trivially holds when $\mathcal{J}_{\boldsymbol{\theta}}=0$.
-However, the collapse of immersion Jacobian results in a degenerate configuration.
-That is, the embedded magnets may have merged or be unable to generate torques.
+It does not affect the workspace volume of the MeSCR, nor any performance metric defined using a left-invariant Riemannian metric.
+Therefore, we let one of the magnets be fixed at the distal end $(L_{k_{N_{m}-1}}=L)$, leaving $N_{m}-1$ decision variables on placements.
+Additionally, we let the flexible segments between adjacent magnets have minimal gaps $(L_{g})$ to mitigate magnetic coupling.
+The orientation of embedded magnets $(\bar{\mathbf{m}}_{k}\in\mathbb{S}^{2})$ is also considered as the design variable.
+We collect the design variables of magnet placements and orientations into the parameter vector $\mathbf{a}\in\mathbb{A}$.
+Under a spatially uniform magnetic field $\mathbf{b}\in\mathbb{B}$ serving as the actuation source, the structural optimization of the MeSCR follows the general form of (9), where the equilibrium mapping $\boldsymbol{\theta}^{\ast}(\mathbf{b};\mathbf{a})$ is implicitly defined by the fixed-point relation in (17).
+Suppose the magnetic field $\mathbf{b}$ satisfies the condition in Theorem 1.
+The manifold $\mathbb{B}$ then constitutes an affine subspace diffeomorphic to $\mathbb{R}^{3}$.
+By Theorem 1, the stiffness matrix $\mathbf{S}$ is nonsingular in a neighborhood of each $\mathbf{b}_{u}\in\mathbb{B}$, ensuring that the equilibrium mapping is differentiable with respect to $\mathbf{b}$.
+Furthermore, by Theorem 2, the Jacobian $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}}=\mathbf{S}^{-1}\mathbf{M}$ satisfies $\operatorname{rk}(\mathbf{S}^{-1}\mathbf{M})=2N_{m}$, implying that $\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}_{u}}=\mathbf{S}^{-1}\mathbf{M}\mathbf{U}_{u}$ is of full rank when $N_{m}\geq 2$.
+Consequently, the immersion $\boldsymbol{\theta}^{\ast}:\mathbb{B}\to\mathbb{T}$ is a smooth mapping, and the structural optimization of the MeSCR can be expressed in the analytical integral form of (12) derived under the smooth manifold assumption.
 Since $\mathcal{J}_{\boldsymbol{\theta}}$ contains complex Hessian terms, we first derive a tractable approximation to facilitate further analysis.
@@
 Suppose the inequality condition in Theorem 1 holds. Then the immersion Jacobian admits the expansion
-$\mathcal{J}_{\boldsymbol{\theta}}^{2}$ $=\det\left(\mathbf{U}_{u}^{\top}\mathbf{M}^{\top}\boldsymbol{\Lambda}^{-2}\mathbf{M}\mathbf{U}_{u}\right)+\epsilon$ (21)
-$\triangleq\tilde{\mathcal{J}}_{\boldsymbol{\theta}}^{2}+\epsilon$
+$\mathcal{J}_{\boldsymbol{\theta}}^{2}=\det\left(\mathbf{U}_{u}^{\top}\mathbf{M}^{\top}\boldsymbol{\Lambda}^{-2}\mathbf{M}\mathbf{U}_{u}\right)+\epsilon\triangleq\tilde{\mathcal{J}}_{\boldsymbol{\theta}}^{2}+\epsilon$ (21)
 where the remainder term $\epsilon$ is bounded above as
@@
 Proof.
-The proof is deferred to the supplementary material. ∎
+It is deferred to the supplementary material. ∎
 It shows that $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$ provides a close approximation of $\mathcal{J}_{\boldsymbol{\theta}}$ in the weak-field regime, with the residual term converging to zero as the magnetic field strength decreases.
-We can now turn to analyze the approximation of the objective $\tilde{\mathcal{Z}}(L_{k})\triangleq\int_{\mathbb{B}}z\ \tilde{\mathcal{J}}_{\boldsymbol{\theta}}\ \Omega_{\mathbb{B}}$.
-Let $\mathbf{M}(\boldsymbol{\theta})$ in (12) be represented as $\mathbf{M}=\mathbf{D}_{k_{m}-1}^{\top}\mathbf{U}_{m}\mathbf{D}_{m}$, we notice
+We can now turn to analyze the approximation of the objective $\tilde{\mathcal{Z}}(\mathbf{a})\triangleq\int_{\mathbb{B}}z\ \tilde{\mathcal{J}}_{\boldsymbol{\theta}}\ \Omega_{\mathbb{B}}$.
+Let $\mathbf{M}(\boldsymbol{\theta})$ in (S15) be represented as $\mathbf{M}=\mathbf{D}_{k_{m}-1}^{\top}\mathbf{U}_{m}\mathbf{D}_{m}$, we notice
 $\mathbf{U}_{m}^{\top}\mathbf{D}_{k_{N_{m}-1}}\boldsymbol{\Lambda}^{-2}\mathbf{D}_{k_{N_{m}-1}}^{\top}\mathbf{U}_{m}$ $=\mathbf{U}_{m}^{\top}\boldsymbol{\Lambda}^{-2}\mathbf{U}_{m}=\boldsymbol{\Lambda}_{m}\!\otimes\!\mathbf{I}_{3}$
@@
 with $\Lambda_{k_{i}}^{m}=\sum_{j=0}^{k_{i}}\frac{\ell_{j}^{2}}{E_{j}^{2}I_{j}^{2}}\approx\frac{1}{k_{0}E^{2}I^{2}}\sum_{j=0}^{i}(L_{k_{j}}-L_{k_{j-1}})^{2}$ as a function of $L_{k}$.
+Lemma 1 also derives the approximation:
+$\mathbf{G}_{\boldsymbol{\theta}}$ $\approx\mathbf{U}_{u}^{\top}\mathbf{D}_{m}^{\top}\left(\boldsymbol{\Lambda}_{m}\otimes\mathbf{I}_{3}\right)\mathbf{D}_{m}\mathbf{U}_{u}$ (22)
+$=\sum_{ij}\Lambda_{ij}^{m}\left(\mathbf{m}_{k_{i}}\mathbf{m}_{k_{j}}^{\top}-\left(\mathbf{m}_{k_{i}}^{\top}\mathbf{m}_{k_{j}}\right)\mathbf{I}_{3}\right).$
 The following lemma shows the relation between the singularity of $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$ and the configuration of $L_{k}$:
@@
 Proof.
-[…]
-$\left(\mathbf{v}^{\top}\otimes\mathbf{I}_{3}\right)\mathbf{U}_{u}\mathbf{D}_{m}=\sum_{i=N_{k}+1}^{N_{m}-1}\left[\mathbf{m}_{k_{i}}\right]_{\times}=\left[\sum_{i=N_{k}+1}^{N_{m}-1}\mathbf{m}_{k_{i}}\right]_{\times}.$
-Consequently, the rank of the expression for $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$ shows that
-$\operatorname{rk}\!\left(\mathbf{U}_{u}^{\top}\mathbf{D}_{m}^{\top}\!\left(\boldsymbol{\Lambda}_{m}\!\otimes\mathbf{I}_{3}\right)\mathbf{U}_{u}\mathbf{D}_{m}\right)\!$
-$=\operatorname{rk}\!\left(\!\left(\mathbf{v}^{\top}\!\!\otimes\mathbf{I}_{3}\right)\!\mathbf{U}_{u}\mathbf{D}_{m}\right)=\operatorname{rk}\left(\left[\sum_{i=N_{k}+1}^{N_{m}-1}\mathbf{m}_{k_{i}}\right]_{\times}\right)\leq 2.$
-That is, the actuation mapping becomes deficient in rank, and we conclude that $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}=0$. ∎
+It is deferred to the supplementary material. ∎
 Lemma 2 indicates that placing all embedded magnets at boundary positions leads to a singularity in $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}$, as expected.
@@
 Theorem 4.
-If the performance index satisfies $z(\mathbf{b}_{u};L_{k})>0$, and the structural parameter $L_{k}$ satisfies the boundary configuration condition in Lemma 4, then the objective $\tilde{\mathcal{Z}}(L_{k})$ attains its minimum at this configuration.
+If the performance index satisfies $z(\mathbf{b}_{u};L_{k})>0$, and the structural parameter $L_{k}$ satisfies the boundary configuration condition in Lemma 2, then the objective $\tilde{\mathcal{Z}}$ attains its minimum at this configuration.
 Proof.
-Since $z>0$ and $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}\geq 0$, it follows that $\tilde{\mathcal{Z}}(L_{k})\geq 0$ for any admissible $L_{k}$.
-When $L_{k}$ satisfies the boundary configuration condition in Lemma 4, we have $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}(\mathbf{b})\equiv 0$ over $\mathbb{B}$, and thus $\tilde{\mathcal{Z}}(L_{k})=0$.
+Since $z>0$ and $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}\geq 0$, it follows that $\tilde{\mathcal{Z}}\geq 0$ for any admissible $L_{k}$.
+When $L_{k}$ satisfies the boundary configuration condition in Lemma 4, we have $\tilde{\mathcal{J}}_{\boldsymbol{\theta}}(\mathbf{b})\equiv 0$ over $\mathbb{B}$, and thus $\tilde{\mathcal{Z}}=0$.
 ∎
-Due to the Jacobian singularity at boundary configurations, the value of the objective becomes insensitive to the gradient of the performance index, thereby weakening the meaningfulness of the solution.
+Due to the Jacobian singularity at boundary configurations, the value of the objective becomes insensitive to the gradient of the performance index.
 However, the minimum spacing constraint prevents the structural parameter $L_{k}$ from reaching such degenerate configurations.
@@
 In this case, the performance gradient becomes the key factor in selecting specific configurations.
-One can also reduce the optimization bias by considering the objective as $\mathcal{Z}/\mathcal{V}$, where $\mathcal{V}\triangleq\int_{\mathbb{B}}\mathcal{J}_{\boldsymbol{\theta}}\Omega_{\mathbb{B}}$ is the total induced volume.
+One can also reduce the optimization bias by considering the objective as $\mathcal{Z}_{\mathrm{vol}}/\mathcal{Z}_{\mathrm{dex}}$.
 This normalized objective reflects the average performance on the configuration space and is insensitive to Jacobian singularity.
-Now we focus on the nontrivial solution of (20). If $z$ is a non-zero performance index, one can further deduce that
-$\operatorname{tr}(\mathbf{G}_{\boldsymbol{\theta}}^{-1}\nabla_{L_{k}}\mathbf{G}_{\boldsymbol{\theta}})=\sum_{i}\frac{\nabla_{L_{k}}\lambda_{i}(\mathbf{G}_{\boldsymbol{\theta}})}{\lambda_{i}(\mathbf{G}_{\boldsymbol{\theta}})}=-2\frac{\nabla_{L_{k}}z}{z}.$ (22)
-Geometrically, this condition implies that optimizing the local performance requires structurally modulating the spectrum of the configuration space metric to counteract its distortion.
-From Lemma 3, we have the following approximation:
-$\mathbf{G}_{\boldsymbol{\theta}}$ $\approx\mathbf{U}_{u}^{\top}\mathbf{D}_{m}^{\top}\left(\boldsymbol{\Lambda}_{m}\otimes\mathbf{I}_{3}\right)\mathbf{D}_{m}\mathbf{U}_{u}$ (23)
-$=\sum_{ij}\Lambda_{ij}^{m}\left(\mathbf{m}_{k_{i}}\mathbf{m}_{k_{j}}^{\top}-\left(\mathbf{m}_{k_{i}}^{\top}\mathbf{m}_{k_{j}}\right)\mathbf{I}_{3}\right)$
-in which $\mathbf{m}_{k}=M(\cos(\theta_{0}^{k})\bar{\mathbf{t}}-\sin(\theta_{0}^{k})\mathbf{v})$.
-While the spectrum is analytically elusive for general $\mathbf{v}\in\mathbb{S}^{2}$, closed-form expressions are still available in the following cases.
+Example: Consider a MeSCR with two axially embedded magnets actuated by a spatially uniform field $\mathbf{b}_{u}$ with $\mathbf{s}^{\top}\bar{\mathbf{t}}=0$, where the global manipulability is to be maximized [21].
+The embedded magnets have identical dipole strength.
+Let one magnet be fixed at the distal end ($L_{k_{1}}=L$), while the position of the other magnet $L_{k_{0}}$ serves as the design variable, subject to the boundary constraint $0<L_{k_{0}}<L$.
+For planar deflection, the joint variable simplifies to $\boldsymbol{\theta}_{i}=\theta_{i}\mathbf{n}$, with $\mathbf{n}\in\mathbb{S}^{2}$ represents the normal vector.
+In this configuration, the MeSCR becomes kinematically equivalent to a planar revolute open-chain mechanism.
+Let $\theta_{i}^{j}\triangleq\sum_{l=i}^{j}\theta_{l}$ denote the cumulative joint angle.
+In the regime of weak magnetic actuation, the global manipulability index has the linearized expression on $\theta_{k_{0}}^{k_{1}}$: $z\approx L_{k_{0}}(L-L_{k_{0}})\left|\sin\theta_{k_{0}}^{k_{1}}\right|$, where the analytical form is derived from the planar 2-joint revolute open-chain mechanism.
+IV-C1 Optimal Orientation
+We parameterize the in-plane magnetic moments by angles $\{\varphi_{k}\}$ with respect to the $\bar{\mathbf{v}}$-axis, where the initial orientation is $\bar{\mathbf{t}}$.
+Let $\varphi_{k_{0}}^{k_{1}}\!\triangleq\!\varphi_{k_{1}}-\varphi_{k_{0}}$ denote the relative orientation, and assume a spatially uniform planar field $\mathbf{b}_{u}\!\in\!\mathbb{R}^{3}$ with support on the $\bar{\mathbf{t}}$-$\bar{\mathbf{u}}$ plane.
+Under the weak-field regime and planar bending, the immersion Jacobian admits the linearized form $\mathcal{J}_{\boldsymbol{\theta}}\;\approx\;c_{0}\big|\sin\!\big(\varphi_{k_{0}}^{k_{1}}+\theta_{k_{0}}^{k_{1}}\big)\big|,$ where $c_{0}>0$ collects geometry- and stiffness-related constants.
+The corresponding global performance functional can be written as
+$\mathcal{Z}\!\left(\varphi_{k_{0}}^{k_{1}}\right)\;=\;c_{1}\!(\sin(\varphi_{k_{0}}^{k_{1}}+2\theta_{k_{0}}^{k_{1}})/4\;-\;\theta_{k_{0}}^{k_{1}}\cos(\varphi_{k_{0}}^{k_{1}})/2),$ (23)
+with $c_{1}>0$ independent of $\varphi_{k_{0}}^{k_{1}}$. In the small-angle limit, the stationarity condition leads to
+$\lim_{\theta_{k_{0}}^{k_{1}}\to 0}\nabla_{\varphi_{k_{0}}^{k_{1}}}\mathcal{Z}=0\;\;\Longrightarrow\;\;\big(\varphi_{k_{0}}^{k_{1}}\big)^{\ast}\;=\;\frac{\pi}{2}\;(\text{mod }\pi),$ (24)
+indicating that the optimal in-plane orientation is orthogonal in the uniform weak field regime.
+Interestingly, it is analogous to the orthogonal joint arrangement in serial manipulators, where perpendicular joint axes maximize manipulability by decoupling motion directions around the nominal configuration.
+Similar orthogonal magnetization has been widely adopted in radially magnetized catheter robots [31, 32], where the magnetic moment is aligned perpendicular to the backbone to enhance steering torque and directional responsiveness under uniform magnetic fields.
+IV-C2 Optimal Placement
+We focus on axially magnetized configurations, as they are the most common design adopted in continuum magnetic robots.
+Applying the approximation from (22), the immersion Jacobian also admits the linearized form:
+$\mathcal{J}_{\boldsymbol{\theta}}\approx$
+[…]
+The cumulative rotation angle can also be approximated using (20) as $\left|\theta_{k_{0}}^{k_{1}}\right|\approx\frac{L-L_{k_{0}}}{EI}\left|\bar{\mathbf{m}}_{k_{1}}^{\top}[\mathbf{n}]_{\times}^{\top}\mathbf{b}_{u}\right|$.
+Substituting this into the above expressions for $z$ and $\mathcal{J}_{\boldsymbol{\theta}}$, we can factor the dependence on $L_{k_{0}}$ out of the integral in the objective.
+The structural optimization problem thus reduces to:
+[…]
+Then, the optimal magnet placement is obtained by using the first-order necessary condition as
+$L_{k_{0}}^{\ast}\approx\begin{cases}1/2.718\ L,\ \text{if }\bar{\mathbf{m}}_{k_{0}}=\bar{\mathbf{m}}_{k_{1}},\\ 2/7\ L,\ \text{if }\bar{\mathbf{m}}_{k_{0}}=-\bar{\mathbf{m}}_{k_{1}}.\end{cases}$
+The result is entirely determined by the structural geometry, independent of material or actuation parameters, and thus exhibits scale invariance.
+Interestingly, when $\bar{\mathbf{m}}_{k_{0}}=\bar{\mathbf{m}}_{k_{1}}$, the ratio of optimal placement $L/L_{k_{0}}^{*}\approx 2.718$ is numerically close to Euler’s number, suggesting a geometric correspondence with the logarithmic spiral structures widely employed in soft continuum design.
 Figure 2: a.) Normalized manipulability and dexterity of MeSCRs with two embedded magnets as functions of normalized magnet position $L_{k_{0}}/L\in[0,1]$: (left) aligned magnetic moments, (right) opposing magnetic moments.
@@
 Color gradient encodes normalized local manipulability.
-Figure 7: a.) Optimization results under $2^{2}$ magnetic orientation combinations.
-The background color maps represent normalized performance values obtained via exhaustive search.
-Black lines indicate trajectories of the gradient-based optimization.
-b.) Optimization results across $2^{3}$ magnetic orientations, comparing the gradient-based method with a heuristic method.
-Bar plots show the configuration at optimized manipulability with convergence curves overlaid.
-III-B1 Analytical Example
-Consider a MeSCR with two axially embedded magnets actuated by a spatially uniform field $\mathbf{b}_{u}\in\Pi$ with $\mathbf{v}^{\top}\bar{\mathbf{t}}=0$, where the global manipulability is to be maximized [23].
-The two magnets have identical dipole strength.
-Let one magnet be fixed at the distal end ($L_{k_{1}}=L$), while the position of the other magnet $L_{k_{0}}$ serves as the design variable, subject to the boundary constraint $0<L_{k_{0}}<L$.
-From Corollary 1, the joint variable simplifies to $\boldsymbol{\theta}_{i}=\theta_{i}\mathbf{n}$.
-In this configuration, the MeSCR becomes kinematically equivalent to a planar $N$-joint revolute open-chain mechanism.
-Let $\theta_{i}^{j}\triangleq\sum_{l=i}^{j}\theta_{l}$ denote the cumulative joint angle.
-In the regime of weak magnetic actuation, the global manipulability index has the linearized expression on $\theta_{k_{0}}^{k_{1}}$: $z\approx L_{k_{0}}(L-L_{k_{0}})\left|\theta_{k_{0}}^{k_{1}}\right|$, where the analytical form is derived from the planar 2-joint revolute open-chain mechanism.
-Similarly, applying the approximate $\mathbf{G}_{m}$ from (23), the immersion Jacobian also admits the linearized form:
-$\mathcal{J}_{\boldsymbol{\theta}}\approx$
-$\begin{cases}\!M^{3}\!\sqrt{\Lambda_{k_{0}}^{m}(\Lambda_{k_{1}}^{m}-\Lambda_{k_{0}}^{m})(3\Lambda_{k_{0}}^{m}+\Lambda_{k_{1}}^{m})}\left|\theta_{k_{0}}^{k_{1}}\!\right|,\text{if }\bar{\mathbf{m}}_{k_{0}}\!=\!\bar{\mathbf{m}}_{k_{1}},\\ \!M^{3}\!\sqrt{\Lambda_{k_{0}}^{m}}(\Lambda_{k_{1}}^{m}-\Lambda_{k_{0}}^{m})\left|\theta_{k_{0}}^{k_{1}}\!\right|,\text{if }\bar{\mathbf{m}}_{k_{0}}=-\bar{\mathbf{m}}_{k_{1}},\end{cases}$
-The cumulative rotation angle can also be approximated using (17) as $\left|\theta_{k_{0}}^{k_{1}}\right|\approx\frac{L-L_{k_{0}}}{EI}\left|\bar{\mathbf{m}}_{k}^{\top}[\mathbf{b}_{u}]_{\times}\mathbf{n}\right|$.
-Substituting this into the above expressions for $z$ and $\mathcal{J}_{\boldsymbol{\theta}}$, we can factor the dependence on $L_{k_{0}}$ out of the integral in the objective.
-The structural optimization problem thus reduces to:
-$\max_{L_{k_{0}}}\begin{cases}\!L_{k_{0}}^{2}(L-L_{k_{0}})^{4}\sqrt{4L_{k_{0}}^{2}\!+\!(L-L_{k_{0}})^{2}},\mbox{if}\ \bar{\mathbf{m}}_{k_{0}}\!=\!\bar{\mathbf{m}}_{k_{1}},\\ \!L_{k_{0}}^{2}(L-L_{k_{0}})^{5},\mbox{if}\ \bar{\mathbf{m}}_{k_{0}}\!=\!-\bar{\mathbf{m}}_{k_{1}}.\end{cases}$
-Then, the optimal magnet placement is obtained by using the first-order necessary condition as
-$L_{k_{0}}^{\ast}\approx\begin{cases}0.368L,\ \text{if }\bar{\mathbf{m}}_{k_{0}}=\bar{\mathbf{m}}_{k_{1}},\\ \frac{2}{7}L,\ \text{if }\bar{\mathbf{m}}_{k_{0}}=-\bar{\mathbf{m}}_{k_{1}}.\end{cases}$
-The result is entirely determined by the structural geometry, independent of material or actuation parameters, and thus exhibits scale invariance.
-Interestingly, when $\bar{\mathbf{m}}_{k_{0}}=\bar{\mathbf{m}}_{k_{1}}$, the ratio of optimal placement $L/L_{k_{0}}^{*}\approx 2.718$ is numerically close to Euler’s number, suggesting a geometric correspondence with the logarithmic spiral structures widely employed in soft continuum design.
-We also evaluated the kinematic dexterity of the MeSCR through minimizing the distortion as defined in [24].
-It is easy to verify that the distortion density satisfies the assumption of Theorem 5.
-Thus, we deduce that the optimal configuration of embedded magnets has a boundary distribution.
-We show numerical simulation results of normalized manipulability and dexterity in Fig.
-6a When the magnetic moments are aligned (left graph), the configuration that maximizes manipulability exhibits good agreement with the theoretical result.
-In contrast, the opposing case (right graph) shows a deviation, which may be attributed to stronger nonlinearity in the index.
-For both cases, the configurations that minimize distortion consistently place the embedded magnet at the distal end, as expected.
-Fig.
-6b illustrates the reachable distal trajectories under varying magnet placements.
-Although the optimal placements yield high manipulability along the path, their spatial extent remains limited.
-This apparent disparity underscores the geometric distinction between local controllability and global reachability, while the latter pertains to the image of the equilibrium map over the actuation submanifold.
-Therefore, any attempt to maximize overall performance must reconcile the trade-off between local dexterity and global coverage, as the two are not generally aligned.
-III-B2 Numerical Examples
-For MeSCRs with three or more embedded magnets, deriving a closed-form solution to the optimization problem becomes increasingly intractable.
-In such cases, we resort to a gradient-based optimization scheme, leveraging the analytical gradient expression provided in (22).
-Fig.
-7a illustrates the optimization trajectories for representative cases with four possible direction combinations.
-The background colormaps depict the globally normalized performance values obtained via exhaustive enumeration.
-Across all cases, the gradient-based method consistently converges to configurations that match the global optima, demonstrating its reliability and consistency with full search results.
-Moreover, the optimized magnet placements for maximizing dexterity closely follow the geometric insights established in Theorem 5.
-In all observed cases, the optimizer implicitly minimizes the net torque induced by the embedded magnets.
-Specifically, when adjacent magnets are aligned, the solution tends to separate them spatially to avoid torque amplification, whereas when they are oppositely oriented, the optimizer brings them closer to enable partial cancellation.
-These outcomes suggest that the dexterity metric inherently penalizes unbalanced torque distributions that restrict the range of achievable actuation directions.
-As shown in Fig.
-7b, the heuristic and gradient-based optimization methods yield consistent results both in terms of optimal configuration and convergence behaviors.
-Combining with the performance maps in Fig.
-7a, it demonstrates that manipulability is highly sensitive to the spatial layout and directions of the embedded magnets.
-In some cases, optimal configurations are observed when magnets are distributed farther apart, highlighting the benefit of spatially diverse actuation sources.
-In comparison to dexterity, manipulability optimization focuses on maximizing the volume of attainable motion directions.
-As a result, the optimizer may favor configurations that introduce local torque in exchange for broader kinematic responsiveness.
-IV Conclusion and Discussion
-This paper presents a unified framework for modeling and optimizing the kinematic performance of multi-magnet embedded soft continuum robots.
-The proposed EPRB model enables a differentiable representation of equilibrium configurations under magnetic actuation, and supports explicit expressions for both the manipulator Jacobian and Hessian in multi-magnet settings.
-The analysis of equilibrium well-posedness yields sufficient conditions for solution uniqueness and continuity under bounded actuation, offering a means to ensure continuous configuration tracking in quasi-static operations.
-The upper bound on controllable degrees of freedom, shown to be twice the number of embedded magnets, offers a simple yet practical rule for relating design parameters to task-space controllability.
-A key contribution of this paper lies in the development of a differential geometric optimization framework, where classical kinematic measures are reparametrized over the actuation space via equilibrium immersions.
-This approach accommodates a wide range of performance criteria, including but not limited to manipulability and dexterity, and admits both analytic and gradient-based solution methods.
-Although demonstrated in the context of MeSCRs, the formulation applies more broadly to continuum structures actuated by distributed torques, and may offer guidance in the design of other soft or compliant mechanisms where actuation and configuration are implicitly coupled.
-Appendix A Proofs of Propositions
-A-A Proof of Proposition 1
-Noticing that
-$\|\mathbf{M}_{i}(\boldsymbol{\theta})\mathbf{b}\|$ $=\left\|\sum_{k\in\mathbb{K},k\geq i}\left(\left[\mathbf{m}_{k}\right]_{\times}^{\top}\mathbf{R}_{0}^{i}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}\right)^{\top}\mathbf{b}(\mathbf{p}_{k})\right\|$
-$\leq\left\|\mathbf{R}_{0}^{i}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{i}}\right\|\sum_{k\in\mathbb{K},k\geq i}\left\|\left[\mathbf{m}_{k}\right]_{\times}\mathbf{b}(\mathbf{p}_{k})\right\|$
-$=\sum_{k\in\mathbb{K},k\geq i}M_{k}B_{k}\triangleq\mathcal{M}_{i},$
-it is convenient to obtain:
-$\|\mathbf{M}(\boldsymbol{\theta})\mathbf{b}\|$ $=\sqrt{\sum_{i=0}^{N-1}\|\mathbf{M}_{i}(\boldsymbol{\theta})\mathbf{b}\|^{2}}\leq\sqrt{\sum_{i=0}^{N-1}\mathcal{M}_{i}^{2}}\triangleq\mathcal{M},$
-and
-$\|\mathbf{M}(\boldsymbol{\theta})\|$ $=\sqrt{\sum_{i=0}^{N-1}\|\mathbf{M}_{i}(\boldsymbol{\theta})\|}\leq\sqrt{N_{m}k_{N_{m}-1}}\max_{k}M_{k}\triangleq\mathcal{M}_{0}.$
-A-B Proof of Proposition 2
-Given that the Hessian matrix is symmetric, it can be expressed equivalently as
-$\frac{\partial^{2}\mathcal{E}_{m}}{\partial\boldsymbol{\theta}\partial\boldsymbol{\theta}^{\top}}=\sum_{k\in\mathbb{K}}\operatorname{sym}\left(\mathbf{D}_{k}^{\top}\mathbf{C}_{k}\mathbf{D}_{k}+\mathbf{Q}^{k}\right)\triangleq\mathbf{S}_{m}.$
-Whence,
-$\mathbf{D}_{k}=\operatorname{blkdiag}\{\mathbf{R}_{0}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{0}},\dots,\mathbf{R}_{0}^{k}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k}},\underbrace{\mathbf{0}_{3\times 3},\dots,\mathbf{0}_{3\times 3}}_{N-k-1}\},$ (24)
-$\mathbf{Q}^{k}=\operatorname{blkdiag}\{\mathbf{Q}_{0},\dots,\mathbf{Q}_{k},\underbrace{\mathbf{0}_{3\times 3},\dots,\mathbf{0}_{3\times 3}}_{N-k-1}\},$
-and
-$\mathbf{C}_{k}=\operatorname{blkdiag}\{2\mathbf{U}_{k\times k}-\mathbf{I}_{k\times k},\underbrace{0\dots 0}_{N-k-1}\}\otimes\left([\mathbf{b}_{k}]_{\times}^{\top}[\mathbf{m}_{k}]_{\times}\right)$
-with $\mathbf{U}_{k\times k}$ denotes the upper triangle matrix with all elements equal to $1$ and $\otimes$ denotes the Kronecker product.
-Then, using the triangle inequality and sub-multiplicative property, the following inequality holds for the spectrum norm of the Hessian matrix:
-$\left\|\mathbf{S}_{m}\right\|$ $=\left\|\sum_{k\in\mathbb{K}}\operatorname{sym}\left\{\mathbf{D}_{k}^{\top}\mathbf{C}_{k}\mathbf{D}_{k}+\mathbf{Q}^{k}\right\}\right\|$ (25)
-$\leq\sum_{k\in\mathbb{K}}\left\|\mathbf{D}_{k}^{\top}\right\|\left\|\mathbf{C}_{k}\right\|\left\|\mathbf{D}_{k}\right\|+\left\|\mathbf{Q}^{k}\right\|$
-$=\sum_{k\in\mathbb{K}}\left\|\mathbf{C}_{k}\right\|+\left\|\mathbf{Q}^{k}\right\|$
-where we used the fact that $\left\|\mathbf{D}_{k}^{\top}\right\|=\left\|\mathbf{D}_{k}\right\|=1$. The spectrum norm of $\mathbf{C}_{k}$ is straightforward to be given by
-$\left\|\mathbf{C}_{k}\right\|$ $=\left\|2\mathbf{U}_{k\times k}-\mathbf{I}_{k\times k}\right\|\left\|[\mathbf{b}_{k}]_{\times}^{\top}[\mathbf{m}_{k}]_{\times}\right\|$ (26)
-$\leq\frac{4k}{\pi}\|\mathbf{b}_{k}\|\|\bar{\mathbf{m}}_{k}\|=\frac{4k}{\pi}M_{k}B_{k}$
-For small rotation angles, we note that
-$\lim_{\boldsymbol{\theta}\to\mathbf{0}}\left\|\mathbf{Q}^{k}\right\|$ $=\lim_{\boldsymbol{\theta}\to\mathbf{0}}\max_{i=0,\dots,k}\left\{\left\|\mathbf{Q}_{i}\right\|\right\}=0.$ (27)
-Combing (25), (26), and (27), the Jacobian of the vector-valued function $\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$ satisfies:
-$\left\|\frac{\partial}{\partial\boldsymbol{\theta}}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}\right\|=\left\|-\mathbf{S}_{m}\right\|\leq\sum_{k\in\mathbb{K}}\frac{4k}{\pi}M_{k}B_{k}\triangleq\mathcal{L}.$
-Therefore, the function is Lipschitz continuous on $\mathbb{R}^{3N}$ with the Lipschitz constant $\mathcal{L}$.
+We also evaluated the kinematic dexterity of the MeSCR by minimizing the distortion index in [22].
+Since the distortion density satisfies the conditions of Theorem 4, the optimal solutions are expected to concentrate near boundary configurations.
+The numerical results in Fig.
+2 confirm this behavior.
+Across both aligned and opposing magnetization modes, the distortion minimizing designs consistently collapse toward the distal end, whereas the manipulability maximizing designs follow the placement trends predicted by the analytical model.
+These outcomes highlight a clear structural organization of the performance landscape.
+Manipulability is governed by the separation and orientation of magnetic sources, while distortion tends to favor extreme placements that suppress curvature variation.
+The reachable trajectories further show that improved local dexterity does not necessarily translate into expanded global coverage.
+This observation reinforces the intrinsic mismatch between local controllability and global reachability and indicates that structural optimization must balance these competing geometric effects rather than optimizing either in isolation.
+IV-D Numerical Results for Multi-magnet Configurations
+For MeSCRs with three embedded magnets, closed-form optimization becomes analytically intractable, thus we employ the gradient-based scheme in Algorithm 1.
+The numerical experiments (see Fig.
+S1) show that the method converges consistently across all orientation patterns and produces solutions that agree with exhaustive or heuristic baselines, which indicates that the objective landscape remains smooth enough for reliable gradient-driven search.
+The structural behaviors observed in the two-magnet case persist in the three-magnet configuration.
+More importantly, the aggregated results reveal clear structural tendencies that persist across all configurations.
+The first two magnets exhibit highly concentrated optimal placements near $L_{k_{0}}^{*}/L\approx 0.18$ and $L_{k_{1}}^{*}/L\approx 0.74$, reflecting a general preference for distributed magnetic sources that enrich the set of attainable torque directions.
+Orientation also plays a decisive role.
+Opposing pairs drive the optimizer to increase their spacing and shift them toward the proximal region, while like-oriented pairs naturally cluster toward one end to exploit larger moment arms.
+These consistent behaviors show that the interaction between magnet placement and orientation organizes the shape of the actuation space and dominates the resulting optimal designs.
+V Optimization under General Fields
+In this section, we extend the proposed structural optimization framework to dipole-generated fields and investigate the resulting equilibrium behavior, kinematic performance, and optimal embedded-magnet configurations.
+V-A Numerical Optimization under Single Dipole Fields
+TABLE I: Geometric and material properties
+Parameters (Description) Value $\mbox{Unit}]$
+Silicone tube $E$ (Young’s modulus) 2020 $\mathrm{MPa}]$
+ODOD (Outside diameter) $1.2$ $\mathrm{mm}]$
+IDID (Inside diameter) $0.8$ $\mathrm{mm}]$
+$L$ (Length) $40$ $\mathrm{mm}]$
+[…]
+[…]
+Micromagnets $B_{r}$ (Remanence) $1.2$ $\mathrm{T}]$
+$L_{m}$ (Length) $5/3/2$ $\mathrm{mm}]$
+$D_{m}$ (Diameter) 11 $\mathrm{mm}]$
+$L_{g}$ (Minimal gap) $5/3/2$ $\mathrm{mm}]$
+[…]
+Actuating Magnet $M_{\mathrm{b}}$ (Dipole moment) $342.86$ $\mathrm{A\cdot m^{2}}]$
+Figure 3: Experimental setup: A permanent magnet attached to a 6-DoF robot arm generates a spatially varying dipole field.
+The MeSCR is placed inside a transparent container, while a downward-facing camera records its deformation for subsequent shape extraction and Jacobian estimation.
+Insets show details of the MeSCR geometry (outer diameter $1.2\text{\,}\mathrm{m}\mathrm{m}$, embedded magnet length $5\text{\,}\mathrm{m}\mathrm{m}$) and an example of the extracted centerline.
+Figure 4: Performance evaluation under three dipole-field task paths.
+a.) Illustration of the three task paths.
+b.)–f.) Normalized kinematic performances with respect to the movable magnet position $L_{k_{0}}/L$ with same and opposite directions of orientation: b.) $\mathcal{Z}_{\mathrm{vol}}$, c.) $\mathcal{Z}_{\mathrm{dis}}$, d.) $\mathcal{Z}_{\kappa}$, e.) $\mathcal{Z}_{\mathrm{vol}}/\mathcal{Z}_{\mathrm{dis}}$, and f.) $\mathcal{Z}_{\mathrm{vol}}/\mathcal{Z}_{\kappa}$.
+Solid curves correspond to numerical results for the three paths, and dashed lines indicate the optimal placements.
+Experimental data are plotted as discrete markers.
+To study the effect of embedded magnet configurations under dipole fields, we evaluate the two principal orientation modes across representative task paths.
+The experimental setup is presented in Fig.
+3, and the results comparison to simulations are summarized in Fig.
+4, Fig.
+S2, and Table S1, which exhibit consistent trends that align with our analytical insights.
+For the same direction magnetization, all kinematic indices preserve their characteristic ordering from the uniform field analysis.
+The increasing actuation strength toward the distal end leads manipulability to prefer distal placement, while indices sensitive to curvature variation attain their maxima at intermediate spacing.
+These behaviors remain stable across task paths and magnet lengths, indicating that constructive torque accumulation dominates the optimization landscape.
+Opposite direction magnetization produces a markedly different pattern.
+The interaction of opposing torques generates a cancellation zone that governs the deformation geometry, causing all performance measures to collapse toward a narrow proximal region.
+The near invariance across paths and magnet lengths confirms that this behavior is determined primarily by the intrinsic structure of the torque field rather than by task-specific sampling.
+The competing tendencies predicted in the uniform field case also persist: increasing magnet separation strengthens net actuation, whereas proximity favors smoother deformation.
+Under dipole actuation, however, the dominant role of cancellation compresses all optimal solutions toward a common proximal configuration.
+These results provide direct experimental validation that magnet orientation is the primary structural factor shaping the Jacobian spectrum and the geometry of the equilibrium manifold.
+Constructive or destructive torque interactions organize the qualitative behavior of all performance measures, while task paths and scaling parameters exert only secondary influence.
+Figure 5: Optimal movable magnet position under varying values of the nondimensional parameter $\nu$. a.) Same direction magnetization. b.) Opposite direction magnetization.
+We further examine the dependence of optimal placement on the nondimensional parameter
+$\nu=\frac{4\pi EI}{N_{m}B_{r}M_{\mathrm{b}}d^{3}AL^{2}}.$
+The results in Fig.
+5 show that for the same direction magnetization, the optimal positions vary smoothly with $\nu$ and retain the established ordering.
+For opposite orientation, the optimal placements exhibit larger variation at small $\nu$ but converge rapidly toward boundary configurations as elasticity increases.
+The presence or absence of internal torque cancellation thus provides the underlying organizing principle for all indices across field strength, task path, and geometry.
+V-B Numerical Optimization under General Fields
+To evaluate whether the dipole–field behaviors observed in the two magnet case generalize to higher dimensional designs, we extend the analysis to MeSCRs with three embedded magnets, allowing two magnets to vary in both position and orientation.
+The experiments cover multiple dipole configurations, task paths, and magnetic strength parameters $(\nu_{1}\!-\!\nu_{5})$, with results summarized in Fig.
+S3.
+A clear and unified set of structural tendencies emerges across all settings.
+The placement hierarchy mirrors that of the two magnet case: constructive torque accumulation drives distal optimality for $\mathcal{Z}_{\mathrm{vol}}$, while indices sensitive to curvature modulation, such as $\mathcal{Z}_{\mathrm{dis}}$ and $\mathcal{Z}_{\kappa}$, consistently favor proximal configurations.
+Ratio based measures converge to stable mid–distal solutions, indicating path invariant behavior even under complex field variations.
+Orientation exhibits the same role identified previously.
+Aligned moments promote torque amplification and favor more concentrated arrangements, whereas partial spreading becomes advantageous when deformation diversity is required.
+The most consistent layouts arise from the combined indices, which naturally balance these competing trends.
+Together with the two magnet dipole–field results, these findings demonstrate that orientation dependent torque interactions and spacing driven actuation diversity form a unified geometric mechanism that persists across increased design dimensionality and external field complexity.
+VI Conclusion
+This work introduced a structural optimization framework that unifies equilibrium mechanics with kinematic performance for magnetically actuated soft continuum robots.
+By formulating magnetic loading as an actuation-dependent equilibrium manifold and evaluating Jacobian spectral indices on its pullback geometry, the approach enables physically consistent performance assessment across both analytical and numerically computed configurations.
+The results across uniform and dipole-generated fields reveal that the attainable motion of magnetic continuum structures is fundamentally organized by the interaction pattern of embedded magnets.
+Constructive torque interactions promote distal or mid–distal arrangements that strengthen net actuation, whereas opposing orientations create intrinsic cancellation regions that compress optimal designs toward proximal configurations.
+These tendencies persist across different actuation paths, field strengths, and magnet counts, indicating that orientation-driven torque geometry is a primary determinant of global kinematic behavior.
+Certain limitations remain.
+The analytical development relies on small-deflection assumptions and a discrete PRB representation, which may reduce accuracy under strong fields or highly compliant materials.
+Magnet–magnet interactions, nonlinear constitutive effects, and dynamic behaviors are not explicitly modeled.
+Extending the approach to richer material descriptions, heterogeneous magnetization, and closed-loop actuation could broaden its practical utility.
+Further characterization of equilibrium-manifold geometry may deepen understanding of reachable sets, while integrating sensing, real-time control, or data-driven design tools could facilitate systematic structure actuation co-design for future magnetic soft continuum robots.